### v2

# Giulia Gaggero and Elisa Gorla

The complexity of solving a random polynomial systemthanks: This research was funded by armasuisse through grant no. CYD-C-2020010.
Giulia Gaggero and Elisa Gorla
1 Introduction
Randomness plays a fundamental role within cryptography. For example, it plays a pivotal role within key generation and cryptographic algorithms are subject to randomness tests. In multivariate cryptography, the public key consists of a multivariate polynomial system and the goal of the attacker is to find a solution of the system. In this context, one typically wishes for the system be as close as possible to, or at least appear as, a random system. A random system is expected to be hard to solve, since the Multivariate Quadratic Problem is not only NP-complete, but also known to be hard to solve on average for a wide range of parameters. From this point of view, e.g., a digital signature scheme whose public keys are sufficiently random is expected to be secure.
In this paper, we discuss what it means for a polynomial system to be random and how hard it is to solve a random polynomial system. In Definition 7 we propose a mathematical formulation for the concept of random system. The definition of randomness that we propose, which we call algebraic randomness, is broad enough to include a vast majority of the systems which are of interest in cryptography. We then specify our definition further in Definition 12. One advantage of this definition is that the property of being random according to Definition 12 can be computationally tested, at least in principle.
In Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32, we prove upper bounds for the degree of regularity and the solving degree of an algebraically random polynomial system, depending on parameters of the system such as the number of equations, the number of variables, and the degrees of the equations. The usefulness of our bounds is twofold: On the one side, our bounds can be used to directly produce bounds on the complexity of computing a Gröbner basis, hence of solving, many systems which are of interest in cryptography. Bounds on the complexity produced in this way have the advantage of being widely applicable and the disadvantage of not always being close to the actual complexity for each system to which they apply. On the other side, our bounds give us an idea of what security one can hope to achieve for a system with given parameters. Therefore, our bounds can be used as a point of comparison for the optimality of a given public key, in the following sense. Say that, in order to forge a signature produced with a given multivariate digital signature scheme, one has to find a solution of a system of $m$ equations of degree $D$ in $n$ variables. Suppose that such a system is algebraically random. Our results provide an upper bound $B$ for the degree of regularity or the solving degree of such a system. Say that one can compute or estimate by a different method the degree of regularity or the solving degree of the specific system and suppose that this turns out to be $C$. Clearly, it must always be that $C\leq B$. However, how far $C$ is from $B$ gives us a measure of how close to optimal the digital signature scheme is, for the given choice of parameters. In fact, since our bounds on the degree of regularity are sharp, there are systems of $m$ equations of degree $D$ in $n$ variables whose degree of regularity is exactly $B$. In other words, if $B$ and $C$ are close, then there is not much space for improvement, since any system with the same parameters as our public key can have degree of regularity or solving degree at most $B$. If on the contrary $B$ and $C$ are far apart, then potentially there is a lot of space for finding a more robust system with the same parameters, since a system with those parameters can have degree of regularity or solving degree up to $B$.
The paper is structured as follows. Section 2 contains some useful preliminaries. In Section 3 we discuss the concept of randomness and propose two definitions of randomness for a polynomial system, in Definition 7 and Definition 12. Throughout the paper, we refer to our notion of randomness as algebraic randomness. We briefly review cryptographic semiregular sequences and semiregular sequences and show that the latter are algebraically random. In Section 4, after recalling some results from commutative algebra, we prove some bounds on the degree of regularity and the solving degree of an algebraically random system consisting of polynomials the same degree. The bounds appear in Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32. Finally, in Section 5 we show how to apply our results to check that the systems associated to GeMSS and Rainbow are not algebraically random.
2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of cardinality $q$ and $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the polynomial ring over $\mathbb{F}_{q}$ in $n$ variables. We start by recalling how a typical multivariate one-way function is constructed. Let $f_{1},\ldots,f_{m}\in R$, and consider the evaluation map
$\begin{array}[]{cccc}\mathcal{F}:&\mathbb{F}_{q}^{n}&\rightarrow&\mathbb{F}_{q}^{m}\\ &\alpha=(\alpha_{1},\ldots,\alpha_{n})&\mapsto&(f_{1}(\alpha_{1},\ldots,\alpha_{n}),\ldots,f_{m}(\alpha_{1},\ldots,\alpha_{n}))\end{array}$
To hide the structure of $\mathcal{F}$, we compose it with two random invertible linear maps $S:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ and $T:\mathbb{F}_{q}^{m}\rightarrow\mathbb{F}_{q}^{m}$. We obtain $\mathcal{P}=T\circ\mathcal{F}\circ S$, a set of $m$ polynomials $p_{1},\ldots,p_{m}$ in $n$ variables over $\mathbb{F}_{q}$. The public key of the multivariate scheme is $\mathcal{P}=(p_{1},\ldots,p_{m})$ and the private key is $\{\mathcal{F},S,T\}$. The trapdoor consists of constructing $\mathcal{F}$ such that $\mathcal{F}^{-1}$ is efficiently computable. Notice that $\mathcal{P}$ should be hard to invert without the knowledge of $S,T$, in particular it should be hard to recover the structure of $\mathcal{F}$ from $\mathcal{P}$.
Finding inverse images with respect to $\mathcal{P}$ corresponds to computing solutions of multivariate polynomial systems. It is well-known that this problem can be solved by computing a Gröbner basis of the system, we refer the interested reader to [13] for more detail. The first algorithm for computing Gröbner bases appeared in the doctoral thesis of Buchberger [11]. Modern algorithms for computing Gröbner bases are based on liner algebra and are more efficient than Buchberger’s. Examples of linear-algebra-based algorithms are $F_{4}$ [20], $F_{5}$ [21], the XL Algorithm [18], MutantXL [12], and their variants. In all of these systems, one computes the reduced row echelon form of the Macaulay matrix associated to the polynomial equations in a given degree, for one or more degrees.
We now describe the main object of linear-algebra-based algorithms [6, Section 1.2]. Fix a term order on $R$ and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subseteq R$ be a system of homogeneous polynomials. The columns of the homogeneous Macaulay matrix $M_{d}$ of $\mathcal{F}$ are labelled by the monomials of $R_{d}$ and arranged in decreasing order. The rows of $M_{d}$ are labelled by polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$. The entry $(i,j)$ of $M_{d}$ is the coefficient of the monomial of column $j$ within the polynomial corresponding to the $i$-th row.
Let now $f_{1},\dots,f_{m}$ be arbitrary (not necessarily homogeneous) polynomials. The columns of the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}$ are labelled by the monomials of $R$ of degree $\leq d$, arranged in decreasing order. The rows of $M_{\leq d}$ correspond to polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})\leq d$. The entries of $M_{\leq d}$ are defined as in the homogeneous case. The rationale behind the use of homogeneous Macaulay matrices for homogeneous systems is that, for a homogeneous system, the Macaulay matrix $M_{\leq d}$ is a block matrix with blocks $M_{d},\ldots,M_{0}$ as blocks.
One computes the reduced row echelon form of the Macaulay matrix, or of its homogeneous version, in one or more degrees. For large enough degree, this produces a reduced Gröbner basis with respect to the chosen order. Some algorithms, as e.g. MutantXL, use a variation called mutant strategy in the non-homogeneous case: If the reduction of the Macaulay matrix $M_{\leq d}$ produces new polynomials $g_{1},\ldots,g_{\ell}$ of degree strictly smaller than $d$, one appends to the reduction of $M_{\leq d}$ the polynomials $m_{i,j}g_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}g_{j})\leq d$, then computes the reduced row echelon form again. Throughout the paper, we refer to the algorithms that employ the mutant strategy as mutant algorithms and to the others as standard algorithms. For a description of the basic version of these algorithms, we refer to [13, Section 3.1] and [14, Section 1].
The computational complexity of computing the reduced row echelon form of the Macaulay matrices $M_{\leq d}$ and $M_{d}$ depends on their size, and hence on the degree $d$. This motivates the next definition.
Definition 1.
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ and let $\tau$ be a term order on $R$. The solving degree of $\mathcal{F}$ with respect to $\tau$ is the least degree $d$ such that Gaussian elimination on the Macaulay matrix $M_{\leq d}$ produces a $\tau$-Gröbner basis of $\mathcal{F}$. We denote by $\operatorname{\mathrm{solv.deg}}_{\tau}^{s}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a standard algorithm and by $\operatorname{\mathrm{solv.deg}}_{\tau}^{m}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a mutant algorithm. When $\tau$ is the degree reverse lexicographic order, we omit the subscript $\tau$.
In general, the solving degree is not invariant under coordinate change. Moreover, it may depend on the algorithm used to perform the Gröbner basis computation. In particular, for mutant algorithms it may be smaller than for standard ones. Finally, the solving degree depends on the choice of a term order on $R$.
The complexity of linear-algebra-based algorithms is dominated by the cost of computing a degree reverse lexicographic Gröbner basis of the system, see [13, Sections 2 and 3] for more detail. Therefore, an upper bound on the solving degree with respect to the reverse lexicographic order yields an upper bound on the complexity of computing a lexicographic Gröbner basis, hence on the complexity of solving the polynomial system.
Let $I$ be a homogeneous ideal of $R$. For an integer $d\geq 0$, we denote by $I_{d}=I\cap R_{d}$ the $\mathbb{F}$-vector space of homogeneous polynomials of degree $d$ in $I$. For $g\in R$ a polynomial, we denote by $g^{\operatorname{\mathrm{top}}}$ the homogeneous part of $g$ of largest degree. E.g., if $g=x^{3}+2xy^{2}-y+1\in\mathbb{F}[x,y]$, then $g^{\operatorname{\mathrm{top}}}=x^{3}+2xy^{2}$. For a polynomial system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$, we denote by $\mathcal{F}^{\operatorname{\mathrm{top}}}\subseteq R$ the homogeneous system $\{f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\}$. Up to doing Gaussian elimination in a matrix whose rows correspond to $f_{1},\ldots,f_{m}$, we may suppose that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent. We will assume this throughout the paper.
The degree of regularity was introduced in [2, 5].
Definition 2.
Let $\mathcal{F}\subseteq R$ be a polynomial system. The degree of regularity of $\mathcal{F}$ is
$d_{\operatorname{reg}}(\mathcal{F})=\left\{\begin{array}[]{cc}\min\{d\geq 0\mid(\mathcal{F}^{\mathrm{top}})_{d}=R_{d}\}&\mbox{ if }\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}\mbox{ for $d\gg 0$}\\ +\infty&\mbox{ otherwise.}\end{array}\right.$
In the cryptographic literature, the degree of regularity is often used as a proxy for the solving degree. This is the case, e.g., in the specification documents of GeMSS [15]. However, this does not always produce reliable estimates. In fact, there are examples in which the gap between the degree of regularity and the solving degree is large, see e.g. [10, Examples 3.2 and 3.3]. A recent result by Semaev and Tenti [27, 28] however shows that, under suitable assumptions, the solving degree of a system with respect to a standard algorithm is at most twice the degree of regularity. Thanks to this result, an upper bound for the degree of regularity yields a proven upper bound for the solving degree.
Theorem 3 ([28, Corollary 3.67] and [27, Theorem 2.1]).
Let $\mathbb{F}=\mathbb{F}_{q}$. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m},x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\subseteq R$ be a polynomial system. If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then for standard algorithms
$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2d_{\operatorname{reg}}(\mathcal{F})-2.$
Notice that, in almost all systems of cryptographic interest, the field size and the degrees of the polynomials are relatively small. Therefore, one expects that Theorem 3 applies to such systems.
Another recent result by Salizzoni [26] shows that the solving degree of a mutant algorithm is at most the degree of regularity plus one, unless the system contains polynomials of large degree.
Theorem 4 ([26, Proposition 3.10]).
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system, then for mutant algorithms
$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{d_{\operatorname{reg}}(\mathcal{F})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}.$
Another algebraic invariant connected to the solving degree of a polynomial system is the Castelnuovo-Mumford regularity. We refer the reader to [13, Section 3.4] and [14] for its definition and a discussion on its relation with the other invariants.
Remark 5.
If $\mathbb{F}=\mathbb{F}_{q}$ and the system $\mathcal{F}$ contains the field equations $x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}$, then $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, therefore the degree of regularity is finite.
Remark 6.
If $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, then
$d_{\operatorname{reg}}(\mathcal{F})=\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}}).$
Therefore:
  • By Theorem 4
    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}$
    for mutant algorithms.
  • If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then by Theorem 3
    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})-2$
    for standard algorithms.
3 Random polynomial systems
Consider a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables. Over a finite field, one often defines a random system as a polynomial system whose coefficients are chosen uniformly at random in the given field. In this paper, we propose a different definition which still captures the intuitive idea of randomness, while allowing us to estimate the degree of regularity of a random system.
Over an infinite field, one may use the concept of genericity from algebraic geometry to define randomness. More precisely, fix a nonempty Zarisky-open subset of $\mathbb{P}^{\binom{n+d_{1}}{n}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}}{n}-1}$, where $\mathbb{P}^{t}$ denotes $t$-dimensional projective space. Define a random system as an element of that open set. This makes sense, since every nonempty Zarisky-open set is dense, hence a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables whose coefficients are chosen uniformly at random is generic with high probability according to this definition. The problem of extending this definition to a finite field is that, over a finite field, a nonempty Zarisky-open set is no longer dense, so the connection with our intuitive idea of randomness is lost. Nevertheless, for a finite field $\mathbb{F}_{q}$, we may define randomness using a Zariski-dense open set defined over the algebraic closure $\overline{\mathbb{F}_{q}}$. While it is not necessarily the case that almost every polynomial system of given degrees with coefficients in $\mathbb{F}_{q}$ is random, this is the case whenever $q$ is large enough, or if we consider a finite extension of $\mathbb{F}_{q}$ of large enough cardinality.
Definition 7.
Let $\mathbb{F}$ be a field and let $\overline{\mathbb{F}}$ be its algebraic closure. Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$. Let $d_{1},\ldots,d_{m}$ be positive integers and let $\mathcal{U}\subseteq\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ be a nonempty Zariski-open set. For a homogeneous polynomial $f$, denote by $[f]\in\mathbb{P}^{\binom{n+\deg(f)-1}{n-1}-1}$ the projective point, whose coordinates are the coefficients of $f$. A homogeneous system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$ for $1\leq i\leq m$ is generic with respect to $\mathcal{U}$ if $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$. It is generic if it is generic with respect to $\mathcal{U}$, for some nonempty Zariski-open set $\mathcal{U}$. An arbitrary system $\mathcal{F}$ is generic or generic with respect to $\mathcal{U}$ if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is.
In the cryptographic literature, random sequences of polynomials are often assumed to be cryptographic semiregular sequences, see e.g. [2, 4]. This is the case in the cryptoanalysis of several systems, as e.g. [15]. The next definition appears in [2, Definition 3.2.1 and Definition 3.2.4][5, Definition 5], and [7, Definition 5 and Definition 9].
Definition 8.
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a homogeneous system.
If $\mathbb{F}\neq\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F})$, one has that $g_{i}\in(f_{1},\ldots,f_{i-1})$.
If $\mathbb{F}=\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2},\ldots,x_{n}^{2}\})$, one has that $g_{i}\in(f_{1},\ldots,f_{i})$.
Arbitrary polynomials $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are a cryptographic semiregular sequence.
In this paper, we use the word cryptographic semiregular sequence in order to distinguish the concept of semiregularity used in the cryptographic literature from the concept of semiregularity originally introduced by Pardue [24, 25], which inspired it. The original definition by Pardue is given over an infinite field $\mathbb{F}$. As we are interested also in dealing with finite fields, we extend it in the natural way.
Definition 9.
Let $\mathbb{F}$ be an infinite field and let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$. Let $I$ be a homogeneous ideal and let $A=R/I$. A polynomial $f\in R_{d}$ is semiregular on $A$ if for every $e\geq d$, the vector space map $A_{e-d}\rightarrow A_{e}$ given by multiplication by $f$ has maximal rank (that is, it is either injective or surjective). If $\mathbb{F}$ is a finite field, let $\overline{R}=\overline{\mathbb{F}}[x_{1},\ldots,x_{n}]$. Then $f$ is semiregular on $A$ if it is semiregular on $\overline{R}/I\overline{R}$.
A sequence of homogeneous polynomials $f_{1},\ldots,f_{m}$ is a semiregular sequence if $f_{i}$ is semiregular on $A/(f_{1},\ldots,f_{i-1})$ for all $1\leq i\leq m$.
It follows from [25, Proposition 1] that, if $\mathbb{F}\neq\mathbb{F}_{2}$, then a semiregular sequence is also a cryptographic semiregular sequence. The converse does not hold, as shown in [25], see the example just below [25, Proposition 1].
The next proposition presents a simple situation in which cryptographic semiregular sequences and semiregular sequences coincide.
Proposition 10.
Let $q>2$ and let $f_{1}=x_{1}^{q}-x_{1},\ldots,f_{n}=x_{n}^{q}-x_{n},f_{n+1}=f\in R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. The sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular if and only if the sequence $f_{1}^{\operatorname{\mathrm{top}}}=x_{1}^{q},\ldots,$ $f_{n}^{\operatorname{\mathrm{top}}}=x_{n}^{q},f_{n+1}^{\operatorname{\mathrm{top}}}=f^{\operatorname{\mathrm{top}}}$ is semiregular.
Proof.
It suffices to show that, if the sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then the sequence $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular. We start by observing that the sequence $x_{1}^{q},\ldots,x_{n}^{q}$ is regular. Therefore, in order to show that $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular, it suffices to show that $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$. If $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular and $d=\deg(f)$, then the Hilbert series of $R/(\mathcal{F}^{\operatorname{\mathrm{top}}})$ is $[(1-z^{d})(1+z+\ldots+z^{q-1})^{n}]$ by [7, Proposition 6]. Here for $p(z)=\sum_{i=0}^{\infty}p_{i}z^{i}\in\mathbb{Z}[[z]]$, we denote by $\delta(p)=\min\{i\geq 0\mid p_{i}\leq 0\}-1$ and define $[p(z)]=\sum_{i=0}^{\delta(p)}p_{i}z^{i}$. Then $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a semiregular sequence by [25, Proposition 1]. ∎
Pardue in [24, 25] shows that Fröberg’s Conjecture [22], a conjecture which has attracted a lot of attention within the commutative algebra community and that is widely believed to hold, is equivalent to the following
Conjecture 11.
Let $\mathbb{F}$ be an infinite field. A generic sequence of polynomials of degrees $d_{1},\ldots,d_{m}$ in $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ is semiregular.
In other words, Fröberg conjectures that the set of semiregular sequences of polynomials of given degrees contains a dense Zarisky-open set. If the conjecture is true, then a sequence of polynomials of given degrees is semiregular with high probability, provided that the ground field has large enough cardinality. It follows that, if $\mathbb{F}=\mathbb{F}_{q}$ with $q\gg 0$ and Fröberg’s Conjecture holds, then a sequence of polynomials of given degrees is a cryptographic semiregular sequence with high probability. In addition, most cryptographic semiregular sequences are also semiregular sequences, as the set of semiregular sequences conjecturally contains a dense open set.
In [2, Section 3.2], Bardet conjectures that a sequence of polynomials with coefficients in $\mathbb{F}_{2}$ is cryptographic semiregular with high probability. This conjecture is motivated by experimental evidence, see also [4, Conjecture 2]. The conjecture was later disproved by Hodges, Molina, and Schlather, who in [23] prove that there are choices of the parameters for which no cryptographic semiregular sequence exists over $\mathbb{F}_{2}$. This is the case, e.g., for $m=1$ and $n>3d_{1}$. In the sequel, we propose a notion of randomness which applies to any choice of the system parameters.
In this paper, we propose two dense Zariski-open sets, which can be used to formalize the intuitive idea of a random system. The set $\mathcal{V}$ corresponds to systems of $m$ polynomials in $n$ variables which contain a regular sequence of length $n$, while the set $\mathcal{U}$ parametrizes systems of $m$ polynomials which contain a regular sequence of $n$ polynomials of the smallest possible degrees. Notice that $\mathcal{V}$ contains $\mathcal{U}$ and, if $m\geq n$, it also contains the set of cryptographic semiregular sequences and that of semiregular sequences.
Definition 12.
Fix $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. For any multiset $\Delta$ of cardinality $n$ contained in the multiset $\{d_{1},\ldots,d_{m}\}$, let $\mathcal{U}_{\Delta}$ be the subset of $\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ whose points correspond to polynomials $f_{1},\ldots,f_{m}\in R$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ contains a regular sequence in the degrees of $\Delta$. Let
$\mathcal{U}=\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}\;\;\;\mbox{ and }\;\;\;\mathcal{V}=\bigcup_{\Delta\subseteq\{d_{1},\ldots,d_{m}\},|\Delta|=n}\mathcal{U}_{\Delta}.$
An algebraically random system of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n$ variables is a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ such that $\deg(f_{i})=d_{i}$ for all $1\leq i\leq m$ and $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$.
It is well-known that $\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}$ is a dense open set for any choice of $1\leq d_{1}\leq\ldots\leq d_{n}$, if $m=n$. This implies that any $\mathcal{U}_{\Delta}$ as in Definition 12 is a dense open set, hence also $\mathcal{U}$ and $\mathcal{V}$ are. Notice that the degree of regularity is finite for all systems in $\mathcal{U}_{\Delta}$ and for any choice of $\Delta$, therefore it is also finite for the systems in $\mathcal{U}$ and $\mathcal{V}$. In particular, the degree of regularity of an algebraically random system is finite. Notice moreover that, if $\mathcal{F}$ is homogeneous, then $\mathcal{V}$ is the set of polynomial systems of degrees $d_{1},\ldots,d_{m}$ for which the degree of regularity is finite.
Notice moreover that, if $m=n$, an algebraically random system is a system for which $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence. Since this case is well-studied in the cryptographic literature, in the sequel we often assume $m>n$.
Remark 13.
For any system $\mathcal{F}$ of equations of degree at least $q$, one has
$\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}.$
Unlike semiregular systems [23], algebraically random systems exist for any choice of the parameters and over every finite field.
Remark 14.
Algebraic random systems over $\mathbb{F}_{q}$ exist for any choice of $n,m$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. This corresponds to the existence of regular sequences of any given degrees in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, as any system of $m$ equations in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ which contains a regular sequence in degrees $d_{1},\ldots,d_{n}$ is algebraically random. Some regular sequences in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{n}$ are for example
$x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},$
where $g_{i}\in\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}]$ and $\deg(g_{i})\leq d_{i}$.
4 The degree of regularity of a random system
In this section, we establish an upper bound for the degree of regularity of an algebraically random system $\mathcal{F}$ consisting of $m$ polynomials of equal degree $D$. In combination with Theorem 3 and Theorem 4, this provides us with an upper bound for the solving degree of a system of algebraically random polynomials of the same degree.
Remark 15.
Notice that $\mathcal{F}\in\mathcal{U}_{\{d,\ldots,d\}}$ for a given $d>0$ if and only if $T\circ\mathcal{F}\circ S\in\mathcal{U}_{\{d,\ldots,d\}}$. In other words, when deciding whether a system of polynomials of equal degree is algebraically random, one can safely ignore the random linear transformations $S$ and $T$ used to disguise the internal system $\mathcal{F}$. This also shows that, for system whose equations are all of the same degree, being an algebraically random system is an intrinsic property of the system and it is not affected by the invertible linear transformations used to disguise the system.
In [10, Section 4], the authors provide an upper bound for the degree of regularity of a system of quadratic polynomials which contains a regular sequence. In this section, we follow the same basic approach and extend it to systems of polynomials of the same degree and systems of polynomials of the same degree to which one adds the field equations. The cases that we treat in this paper are technically more challenging and require the use of more sophisticated results from commutative algebra.
We start by introducing the family of lex-segment ideals. A conjecture by Eisenbud, Green, and Harris will allow us to reduce to these ideals, when estimating the regularity of ideals generated by algebraically random systems. Throughout the section, we fix the lexicographic term order on $R$ with $x_{1}>x_{2}>\ldots>x_{n}$.
Definition 16.
A monomial ideal $I\subseteq R$ is a lex-segment ideal if it has the property that if $u,v\in R$ are monomials of the same degree such that $u\geq_{\operatorname{lex}}v$ and $v\in I$, then $u\in I$.
Let $C$ and $c_{1}\leq\ldots\leq c_{n}$ be non negative integers. An ideal $\mathcal{L}\subseteq R$ is a $(c_{1},\ldots,c_{n};C)$-LexPlusPowers (LPP) ideal if $\mathcal{L}=(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})+L$, where $L$ is a lex-segment ideal generated in degree $C$.
Notation 17.
Let $I\subseteq R$ be a homogeneous ideal containing a regular sequence of polynomials of degrees $c_{1}\leq\ldots\leq c_{n}$. For each $C\geq 0$, we denote by $\operatorname{LPP}(I;c_{1},\ldots,c_{n};C)$ the $(c_{1},\ldots,c_{n};C)$-LPP ideal $\mathcal{L}=(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})+L$ such that $\dim(I_{C})=\dim(\mathcal{L}_{C})$. We make $L$ unique by choosing the largest lex-segment ideal generated in degree $C$ for which the equality $\mathcal{L}=(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})+L$ holds.
Example 18.
Let $I\subseteq\mathbb{F}[x,y,z]$ be generated by a homogeneous regular sequence of polynomials of degrees $1,3,4$. The $(1,3,4;3)$-LPP ideal $\mathcal{L}$ such that $\dim(I_{3})=7=\dim(\mathcal{L}_{3})$ is $\mathcal{L}=(x,y^{3},z^{4})$.
For any lex-segment ideal $L$ generated in degree $3$ with $\dim(L_{3})\leq 7$, one has $L\subseteq\mathcal{L}$. Hence one may write
$\mathcal{L}=(x,y^{3},z^{4})+L$
as in Definition 16 and choose e.g. $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2})$, or $L=(x^{3})$. However, Notation 17 prescribes that we choose the largest $L$ with respect to containment, i.e.
$L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2},y^{3})$
is the lex-segment ideal generated by the first $7$ cubic monomials in the lexicographic order.
The next conjecture appears as [19, Conjecture ($V_{m}$)]. It has been settled in several cases and it is widely believed to hold within the commutative algebra community. For an introduction to the conjecture and an excellent survey of known cases, we refer the interested reader to [17]. Here we state it in a weak form, which is what we need in the sequel.
Conjecture 19 (Eisenbud-Green-Harris Conjecture).
Let $I\subseteq R$ be a homogeneous ideal containing a regular sequence of polynomials of degrees $c_{1}\leq\ldots\leq c_{n}$. Then
$\operatorname{reg}(I)\leq\operatorname{reg}(\operatorname{LPP}(I;c_{1},\ldots,c_{n};C))$
for all $C\geq c_{n}$.
In order to estimate the degree of regularity of our systems, we use the following result by Caviglia and De Stefani.
Proposition 20 ([16, Lemma 2.3]).
Let $c_{1}\leq\ldots\leq c_{n}$ and $2\leq C\leq\sum_{i=1}^{n}(c_{i}-1)$. Let $\mathcal{L}=(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})+L$ be a $(c_{1},\ldots,c_{n};C)$-LPP ideal, and assume that $\mathcal{L}\neq(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})$. Let $u=x_{k}^{t_{k}}v$, with $t_{k}\neq 0$ and $v\in\mathbb{F}[x_{k+1},\ldots,x_{n}]$, be the smallest monomial with respect to the lexicographic order which belongs to $L$ and has degree $C$. Then
$\operatorname{reg}(\mathcal{L})=t_{k}+\sum_{i=k+1}^{n}(c_{i}-1).$
Our first result is an explicit bound for the degree of regularity of an algebraically random system of polynomials of the same degree. In order to make the proof more readable, we introduce the following notation.
Notation 21.
If $u\in R_{D}$ is a monomial that only involves the variables $x_{k},\ldots,x_{n}$ and has degree $a$ in $x_{k}$, we say that $u$ is a $(D,k,a)$-type monomial.
In the next theorem, we provide an explicit formula for the degree of regularity of an LPP ideal with given parameters. Thanks to Conjecture 19, this yields an upper bound for the degree of regularity of an algebraically random system of polynomials of the same degree. In the statement of the theorem, $\sigma_{k,t}$ is the position of the smallest $(D,k,D-t)$-type monomial in the ordered list of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$, sorted in decreasing lexicographic order. In particular, $\sigma_{1,0}=0$ and $\sigma_{n-1,D-1}$ is the number of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$.
Theorem 22.
Assume that Conjecture 19 holds. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system and assume without loss of generality that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
$\sigma_{k,t}=\sum_{i=1}^{k}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}-\sum_{j=t+1}^{D-1}\binom{n-k-1+j}{j}.$
If $\mathcal{F}$ is an algebraically random polynomial system, then
$d_{\operatorname{reg}}(\mathcal{F})\leq(D-t)+(n-k)(D-1).$
Proof.
If $m=n$, then $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence of $n$ polynomials of degree $D$, hence
$d_{\operatorname{reg}}(\mathcal{F})=n(D-1)+1.$
Suppose therefore that $m>n$ and let $J$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}$. Consider the lexicographic order on $R$ and let
$\mathcal{L}=\operatorname{LPP}(J;D,\ldots,D;D)=(x_{1}^{D},\ldots,x_{n}^{D})+L,$
be the LPP ideal with $L$ the largest lex-segment ideal generated in degree $D$ such that $\dim(\mathcal{L}_{D})=m.$ Since both $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ and $x_{1}^{D},\ldots,x_{n}^{D}$ are linearly independent, we have
$\dim\left(\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle\right)=m-n.$
For $11\leq k\leq n-$ 1 and $10\leq t\leq D-$ 1, the number of $(D,k,D-t)$-type monomials is
$\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]_{t})=\binom{n-k-1+t}{t},$
hence
$\begin{split}\sigma_{k,t}&=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}+\sum_{j=1}^{t}\binom{n-k-1+j}{j}\\ &=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\dim(\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}])_{j}+\sum_{j=1}^{t}\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}])_{j}\end{split}$
is the number of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ different from $x_{1}^{D},\ldots,x_{n}^{D}$ and bigger than or equal to $x_{k}^{D-t}x_{n}^{t}$, the smallest $(D,k,D-t)$-type monomial. In other words, the monomial $x_{k}^{D-t}x_{n}^{t}$ is in position $\sigma_{k,t}$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. Notice moreover that $\sigma_{1,0}=0$ and $\sigma_{k,0}=\sigma_{k-1,D-1}$ for $12\leq k\leq n-$ 1.
If $u$ is the smallest monomial in $\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle$, then $u$ is in position $m-n$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. If $\sigma_{k,t-1}<m-n\leq\sigma_{k,t}$ for some $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1, then $u$ is a $(D,k,D-t)$-type monomial. Since $0=\sigma_{1,0}<m-n\leq\dim(\mathbb{F}_{q}[x_{1},\ldots,x_{n}]_{D})/(x_{1}^{D},\ldots,x_{n}^{D})_{D}=\sigma_{n-1,D-1}$, then $m-n$ always belong to one of the intervals above.
The ideal $L$ is generated by the degree $D$ monomials which are greater than or equal to $u$, unless the monomial following $u$ in lexicographic decreasing order is a pure power. In that case, $u=x_{k}x_{n}^{D-1}$ for some $11\leq k\leq n-$ 1, i.e. $m-n=\sigma_{k,D-1}$, and the smallest degree $D$ monomial in $L$ is $x_{k+1}^{D}$. Then $L$ is generated by the degree $D$ monomials which are greater than or equal to $x_{k+1}^{D}$, which is a $(D,k+1,D)$-type monomial. In both situations
$\operatorname{reg}(\mathcal{L})=(D-t)+(n-k)(D-1)$
by Proposition 20. The thesis now follows from observing that
$d_{\operatorname{reg}}(\mathcal{F})=\operatorname{reg}(J)\leq\operatorname{reg}(\mathcal{L}),$ (1)
where the inequality follows from Conjecture 19. ∎
Example 23.
In this example we show how to compute the bound from Theorem 22 for concrete choices of the parameters. Let $n=6$ and $D=3$, so =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{6}]_{3}$. Then $1\leq k\leq 5$ and $0\leq t\leq 2$. The values of $\sigma_{k,t}$ are
$\sigma_{k,t}$ $t=0$ $t=1$ $t=2$
$k=1$ 0 5 20
$k=2$ 20 24 34
$k=3$ 34 37 43
$k=4$ 43 45 48
$k=5$ 48 49 50
If $m=12$, then $m-n=6$ and $\sigma_{1,1}<6\leq\sigma_{1,2}$. Hence $k=1$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 11.$
If $m=42$, then $m-n=36$ and $\sigma_{3,0}<36\leq\sigma_{3,1}$. Hence $k=3$, $t=1$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 8.$
If $m=54$, then $m-n=48$ and $\sigma_{4,1}<48\leq\sigma_{4,2}$. Hence $k=4$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 5.$
Remark 24.
The upper bound of Theorem 22 is decreasing as a function of $m$, as one would expect. In particular, as $m-n$ passes from an interval $(\sigma_{k,t-1},\sigma_{k,t}]$ to the next, the upper bound decreases by one. The largest value of the bound is obtained in the case $m=n$, which corresponds to $\mathcal{F}^{\operatorname{\mathrm{top}}}$ being a regular sequence. In this case, the value for the bound is well-know and is $n(D-1)+1$. The smallest value for the bound is obtained in the case $m-n=\sigma_{n-1,D-1}$, which corresponds to $\langle\mathcal{F}^{\operatorname{\mathrm{top}}}\rangle+\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle=R_{D}$. In this case, the value for the bound is easily seen to be $D$.
Remark 25.
The upper bound produced in Theorem 22 is sharp for all values of $m,n,D$. In fact, it is met by any system $\mathcal{F}$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ is a $(D,\ldots,D;D)$-LPP ideal.
Combining Theorem 22 and Theorem 4, we obtain the following.
Corollary 26.
Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. If Conjecture 19 holds, then
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F})\leq D-t+1+(n-k)(D-1).$
Proof.
The upper bound found in Theorem 22 for the degree of regularity of $\mathcal{F}$ is bigger than or equal to $D$ for all $k$ and $t$. The thesis then follows from Theorem 4. ∎
We now wish to apply Theorem 22 to a system which contains the field equations.
Remark 27.
After reducing the equations of $\mathcal{F}$ modulo the field equations, they have degree at most $q-1$ in each variable, hence total degree at most $n(q-1)$. Therefore, when adding the field equations to a polynomial system of degree $D$, we may always assume that
$D\leq n(q-1).$
Combining Theorem 22 and Theorem 3, one obtains the following. By Remark 27, we may assume without loss of generality that $D\leq n(q-1)$.
Corollary 28.
Assume that Conjecture 19 holds. Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials, with $D\leq n(q-1)$. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. If $d_{\operatorname{reg}}(\mathcal{F})\geq q$, then
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2\min\{n(q-1),(n-k+1)(D-1)-t\}.$
Proof.
First, since $(x_{1}^{q},\ldots,x_{n}^{q})_{n(q-1)+1}=R_{n(q-1)+1}$, one has
$d_{\operatorname{reg}}(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})\leq n(q-1)+1.$
If $D<q$, then
$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{q-1}=(\mathcal{F}^{\operatorname{\mathrm{top}}})_{q-1}\neq R_{q-1},$
since $d_{\operatorname{reg}}(\mathcal{F})\geq q$ by assumption. Therefore, $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq q=\max\{D,q\}$. If $q\leq D\leq n(q-1)$, then
$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1},$
hence $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq D=\max\{D,q\}$. This shows that the assumptions of Theorem 3 are satisfied. The thesis now follows by combining Theorem 22, Theorem 3, and the observation that
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq d_{\operatorname{reg}}(\mathcal{F}).$
Remark 29.
While the estimates of Corollary 26 and of Corollary 28 hold for every $D$, they are most relevant for $D\leq q$. In fact, for $D>q$ we obtain tighter upper bounds on the degree of regularity - hence on the solving degree - of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in Theorem 34 by inspecting the degree $D$ part of the system $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}$. This corresponds to the fact that, whenever the degree of the equations of the system is larger than or comparable to the field size, it is convenient to add the field equations to the system before computing a Gröbner basis.
Next we estimate the degree of regularity and the solving degree of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in the case when $D\geq q$. By Remark 27, we may assume that $D\leq n(q-1)$.
First, we observe that one can easily derive a lower bound on the degree of regularity.
Remark 30.
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $q\leq D\leq n(q-1)$. Since $(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1}$, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq D.$
We now want to derive an upper bound. We start by computing the number of linearly independent homogeneous polynomials of degree $D$ as a function of $n,D,q$.
Remark 31.
Assume that $q\leq D\leq n(q-1)$. A standard Hilbert function computation shows that the number of homogeneous polynomials in $n$ variables of degree $D$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is
$\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}=\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}.$
It follows that the assumption that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ holds on a dense open set, whenever
$m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}.$ (2)
If inequality (2) is not satisfied, then $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ cannot be linearly independent.
The simplest case to treat is that of very overdetermined systems, more specifically the case when $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are too many to be linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$. This happens when $m$ is larger than the bound from Remark 31. The next proposition shows that, in such a situation, the degree of regularity is equal to the degree of the equations of the system.
Proposition 32.
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, where $q\leq D\leq n(q-1)$. If $m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}$, then there is a dense open set $\mathcal{W}$ such that, if $\mathcal{F}\in\mathcal{W}$, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})=D.$
Moreover,
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2D-2$
and
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq D+1.$
Proof.
Since by assumption
$m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}=\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D},$
then there is an open set $\mathcal{W}$ of $m$-tuples of polynomials of degree $D$ such that
$\langle f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\rangle+(x_{1}^{q},\ldots,x_{n}^{q})_{D}=R_{D}.$
Since $(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1}$, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})=D.$
The rest of the statement now follows from Theorem 3 and Theorem 4. ∎
The next theorem yields an upper bound on the degree of regularity of an algebraically random system of equations of degree larger than the field size to which we add the field equations. For such a system, the bound is tighter than the one from Corollary 28. We start with a preparatory lemma, whose proof follows directly from the definition.
Lemma 33.
Let $u,v$ be monomials of type $(D,k,a)$ and $(D,h,b)$, respectively. If $u\geq v$, then either $k<h$ or $k=h$ and $a\geq b$. In particular, if $u,v,w$ are monomials such that $u\geq v\geq w$ and $u$ and $w$ have the same type, then $v$ also has the same type as $u$ and $w$.
Theorem 34.
Assume that Conjecture 19 holds. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, with $q\leq D\leq n(q-1)$. Assume that $m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$ and that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$. Let $11\leq t\leq q-$ 1 and $11\leq k\leq n-$ 1 be such that $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
and
$\eta_{k,t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1}.$
Let
$B=q-t+(n-k)(q-1).$
If $m=\sigma_{k,t}$ and either $t\neq q-1$ and $D\geq 2q-t-1$, or $t=q-1$ and $D\geq 2q-1$, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B-1,$
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2(B-2),$
and
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B+1.$
In any other case
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B,$
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2(B-1),$
and
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B+1.$
Proof.
We start by proving that $\eta_{k,t}$ is the number of $(D,k,q-t)$-type monomials in $R$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$, that is
$\eta_{k,t}=\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}.$ (3)
Since $x_{k+1}^{q},\ldots,x_{n}^{q}$ is a regular sequence in $\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]$, a standard computation involving Hilbert series yields the explicit formula
$\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1},$
where we notice that $\eta_{k,t}\neq 0$ only if $D+t-q\leq(n-k)(q-1)$. This establishes the equality in (3). Similarly to the proof of Theorem 22, we notice that
$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
is the number of monomials of degree $D$ which do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$ and are greater than or equal to $x_{k}^{q-t}x_{n}^{D-q+t}$, the smallest $(D,k,q-t)$-type monomial. Moreover, $\sigma_{k-1,q-1}=\sigma_{k,0}$ for $12\leq k\leq n-$ 1.
Let $I\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\}$. Let $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$ be the $(q,\ldots,q;D)$-LPP ideal such that
$\dim(I_{D})=\dim(\mathcal{L}_{D})=m+\dim(x_{1}^{q},\ldots,x_{n}^{q})_{D}.$
Hence $\mathcal{L}$ is minimally generated by $x_{1}^{q},\ldots,x_{n}^{q}$ and $m$ monomials of degree $D$ that do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$. Recall that, by assumption, $L$ is the largest lex-segment ideal generated in degree $D$ such that $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$, i.e., $L_{D}\supseteq(x_{1}^{q},\ldots,x_{n}^{q})_{D}$.
Let $u$ be the smallest degree $D$ monomial in $\mathcal{L}/(x_{1}^{q},\ldots,x_{n}^{q})$. Notice that $u$ is a $(D,k,q-t)$-type monomial, since $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. Let $v$ be the smallest degree $D$ monomial in $L$. If $u$ is the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$, then $u=x_{k}^{q-t}x_{k+1}^{q-1}\cdots x_{n}^{q-1}$ and $D=q-t+(n-k)(q-1)$. Moreover, $\mathcal{L}_{D}=R_{D}$ and $v=x_{n}^{D}$ is a $(D,n,D)$-type monomial.
Assume now that $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ and let $w$ be the monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ which follows $u$ in decreasing lexicographic order. Then $u\geq v\geq w$ and $v$ is the degree $D$ monomial next to $w$ in increasing lexicographic order. If $m\neq\sigma_{k,t}$, then $w$ has type $(D,k,q-t)$, hence so does $v$ by Lemma 33. Suppose therefore that $m=\sigma_{k,t}$. Write $D-q+t=(n-\ell)(q-1)+r$, where $10\leq r<q-$ 1. Notice that $D-q+t<(n-k)(q-1)$, since $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$. Then $\ell>k$. In this situation, $u=x_{k}^{q-t}x_{\ell}^{r}x_{\ell+1}^{q-1}\cdots x_{n}^{q-1}$ and $w=x_{k}^{q-t-1}x_{k+1}^{q-1}\cdots x_{k+n-\ell}^{q-1}x_{k+n-\ell+1}^{r+1}$. Notice that $w$ is a $(D,k,q-t-1)$-type monomial if $t\neq q-1$ and a $(D,k+1,q-1)$-type monomial if $t=q-1$. If $w$ is not the smallest degree $D$ monomial of its type in $R_{D}$, then $v$ has the same type as $w$. If $w$ is the smallest degree $D$ monomial of its type in $R_{D}$, then $w=x_{k}^{q-t-1}x_{k+1}^{D-q+t+1}$ in the case $t\neq q-1$ and $w=x_{k+1}^{q-1}x_{k+2}^{D-q+1}$ in the case $t=q-1$. Since $w\not\in(x_{1}^{q},\ldots,x_{n}^{q})$, this is only possible if $D\leq 2q-t-2$ for $t\neq q-1$ and $D\leq 2q-2$ for $t=q-1$. In this situation, $v$ has type $(D,k,q-t)$ in the case $t\neq q-1$ and type $(D,k+1,q)$ in the case $t=q-1$.
If Conjecture 19 holds, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq\operatorname{reg}(\mathcal{L}).$ (4)
Moreover, by Proposition 20
$\operatorname{reg}(\mathcal{L})=\left\{\begin{array}[]{ll}q-t-1+(n-k)(q-1)&\mbox{if }m=\sigma_{k,t}\mbox{ and }\\ &\mbox{either }t\neq q-1\mbox{ and }D\geq 2q-t-1,\\ &\mbox{or }t=q-1\mbox{ and }D\geq 2q-1,\\ \\ q-t+(n-k)(q-1)&\mbox{else.}\end{array}\right.$ (5)
The bound on the degree or regularity now follows from (4) and (5). The bounds on the solving degree follow from the bound on the degree of regularity, Theorem 3, and Theorem 4. ∎
5 Applications to the study of GeMSS and Rainbow
GeMSS and Rainbow were the only multivariate schemes in Round 3 of the NIST Post-Quantum Cryptography Standardization process. They are based on modifications of HFE (Hidden Field Equation) and UOV (Unbalanced Oil and Vinegar), respectively. They were subsequently revealed to be insecure and susceptible to a MinRank Attack, see e.g. [3, 8, 9, 1], therefore they were excluded from the NIST competition.
In this section, we demonstrate how to use our results to show that these systems are far from being algebraically random. We do so by computing their degree of regularity or solving degree for small instances and comparing it with the upper bounds that we obtained in Theorem 22 and Theorem 34. The degree of regularity and solving degree of the systems associated to GeMSS and Rainbow are much smaller than the corresponding invariants for an algebraically random system with the same parameters, which reveals the presence of a hidden structure that may be used to mount an ad-hoc attack, as it was done in practice with the MinRank Attacks mentioned above.
We used Magma to compute the solving degree and Singular to compute the degree of regularity. The values that we obtain (and that we indicate in the tables below) are almost always the same for systems with the same parameters. For each choice of the parameters in the table, we produce ten instances of the public key $\mathcal{PK}=\{p_{1},\ldots,p_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of the chosen scheme. For each one of them, we choose a random vector $s=(s_{1},\ldots,s_{m})\in\mathbb{F}_{q}^{m}$ as a signature (in case the chosen vector is not a valid signature, we replace it with another randomly chosen vector). In order to forge the signature, an attacker may want to solve the system $\mathcal{PK}s=\{p_{1}-s_{1},\ldots,p_{m}-s_{m}\}$. We make the system $\mathcal{PK}s$ square by assigning random values to the last $n-m$ variables. This yields a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$.
Since in GeMSS we work over $\mathbb{F}_{2}$, we add to the system $\mathcal{F}$ the field equations =$\mathcal{E}=\{x_{1}^{2}+x_{1},\ldots,x_{m}^{2}+x_{m}\}$. In the next table we compare the experimental results we obtained for GeMSS with the bounds from Theorem 34. The experiments show that both the solving degree and the degree of regularity of $\mathcal{F}$ can be more than twice the solving degree of $\mathcal{F}\cup\mathcal{E}$. This confirms the intuition that adding the field equations is a good strategy in order to solve the system $\mathcal{F}$ over $\mathbb{F}_{2}$.
Unfortunately we were able to compute the degree of regularity of $\mathcal{F}\cup\mathcal{E}$ only for small values of the parameters. In the next table the first three columns contain the parameters of the cryptosystem, and the fourth the number of polynomials and variables that appear in $\mathcal{F}$. The columns labelled '$d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular. The columns labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' and '$\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' are the bounds given by the Theorem 34 for the chosen parameters and a standard Gröbner basis algorithm.
n, D a v m $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
12, 4 1 1 11 5 10 3 18
8, 9 1 1 7 3 6 3 10
8, 9 1 2 7 3 6 3 10
8, 9 2 1 6 3 5 3 8
8, 9 2 2 6 3 5 3 8
24, 4 1 1 23 22 4 42
24, 4 1 2 23 22 4 42
24, 4 1 3 23 22 4 42
24, 4 2 1 22 21 4 40
24, 4 2 2 22 21 4 40
24, 4 3 1 21 20 4 38
While GeMSS is random according to Definition 12, the experimental results make it clear that both the degree or regularity and the solving degree of GeMSS are far from the largest values that one can find for a system of those parameters according to Theorem 34. This indicates that, for the same parameters, one should be able to find systems for which the complexity of computing a Gröbner basis is much larger. More importantly, it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the MinRank attacks mentioned in the opening paragraph).
For Rainbow, we choose to work over $\mathbb{F}_{4}$ and $\mathbb{F}_{9}$. Since $\mathcal{F}$ is a square system, $\mathcal{F}$ is algebraically random if and only if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence. This turns out to be the case in most of the examples that we computed and in that case
$d_{\operatorname{reg}}(\mathcal{F})=m+1$ (6)
by Theorem 22. This is confirmed by our computations.
Since the systems coming from this scheme are quadratic, adding the field equations may increase the solving degree of the system. However, for the small values of $q$ that we tried in our experiments, we find that in all cases but one the solving degree decreases when adding the field equations. This makes sense, as the degree $q$ of the equations that we add is never larger than the solving degree of the system to which we add them. In the next table we summarize the results that we obtained in our computational experiments. Since in our experiments the solving degree of $\mathcal{F}\cup\mathcal{E}$ is almost always smaller than that of $\mathcal{F}$, in our examples $\mathcal{F}\cup\mathcal{E}$ is the relevant system to consider, that is, the system that one wants to try to solve. Therefore, we consider the degree of regularity and solving degree of $\mathcal{F}\cup\mathcal{E}$. In our table, we compare the degree of regularity and the solving degree of $\mathcal{F}\cup\mathcal{E}$ with the upper bounds from Theorem 22 and Corollary 28. We use Corollary 28 as the algorithm implemented in Magma is a standard one. The first three columns contain the chosen values for the parameters and the number of polynomials and variables that appear in $\mathcal{F}$. The columns labelled '$d_{\operatorname{reg}}(\mathcal{F})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular. The column labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contains the bounds from Corollary 28. We do not include the values of the degree of regularity and solving degree of $\mathcal{F}$ in the table, as experimentally we find that the system $\mathcal{F}\cup\mathcal{E}$ can always be solved more efficiently than the system $\mathcal{F}$.
$q$ $[v_{1},o_{1},o_{2}]$ $m$ $d_{\operatorname{reg}}(\mathcal{F})$ $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
4 $[3,2,2]$ 4 4/5 4 4 8
4 $[3,3,3]$ 6 6/7 5 5 12
4 $[7,5,5]$ 10 10/11 6 6 20
9 $[3,2,2]$ 4 5 5 9
9 $[7,5,5]$ 10 11 9/10 10 20
Notice that Corollary 28 does not apply to the case $q=9$ and $[v_{1},o_{1},o_{2}]=[3,2,2]$, since $d_{\operatorname{reg}}(\mathcal{F})=5$. For these parameters, the bound from Corollary 28 would yield $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})\leq 8$. However, the bound does not hold in this case, as our experiments show.
As for GeMSS, we observe that the values that we computed for the solving degree of $\mathcal{F}\cup\mathcal{E}$ are far from the upper bounds predicted by Corollary 28. We conclude that, also in this case, one expects to find systems with the same parameters as these instances of Rainbow and for which the complexity of computing a Gröbner basis is larger. More importantly, once again it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the attacks that we mentioned at the start of the section).

### v3

@@
 Giulia Gaggero and Elisa Gorla
-The complexity of solving a random polynomial systemthanks: This research was funded by armasuisse through grant no. CYD-C-2020010.
-Giulia Gaggero and Elisa Gorla
+Abstract
+Many systems of interest in cryptography consist of equations of the same degree.
+Under the assumption that the degree of regularity is finite, we prove upper bounds on the degree of regularity of a system of equations of the same degree, with or without adding the field equations to the system.
+The bounds translate into upper bounds on the solving degree of the systems, and hence on the complexity of solving them via Gröbner bases methods.
+Our bounds depend on the number of equations in the system, the number of variables, and the degree of the equations.
 1 Introduction
-Randomness plays a fundamental role within cryptography.
-For example, it plays a pivotal role within key generation and cryptographic algorithms are subject to randomness tests.
-In multivariate cryptography, the public key consists of a multivariate polynomial system and the goal of the attacker is to find a solution of the system.
-In this context, one typically wishes for the system be as close as possible to, or at least appear as, a random system.
-A random system is expected to be hard to solve, since the Multivariate Quadratic Problem is not only NP-complete, but also known to be hard to solve on average for a wide range of parameters.
-From this point of view, e.g., a digital signature scheme whose public keys are sufficiently random is expected to be secure.
-In this paper, we discuss what it means for a polynomial system to be random and how hard it is to solve a random polynomial system.
-In Definition 7 we propose a mathematical formulation for the concept of random system.
-The definition of randomness that we propose, which we call algebraic randomness, is broad enough to include a vast majority of the systems which are of interest in cryptography.
-We then specify our definition further in Definition 12.
-One advantage of this definition is that the property of being random according to Definition 12 can be computationally tested, at least in principle.
-In Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32, we prove upper bounds for the degree of regularity and the solving degree of an algebraically random polynomial system, depending on parameters of the system such as the number of equations, the number of variables, and the degrees of the equations.
-The usefulness of our bounds is twofold: On the one side, our bounds can be used to directly produce bounds on the complexity of computing a Gröbner basis, hence of solving, many systems which are of interest in cryptography.
-Bounds on the complexity produced in this way have the advantage of being widely applicable and the disadvantage of not always being close to the actual complexity for each system to which they apply.
-On the other side, our bounds give us an idea of what security one can hope to achieve for a system with given parameters.
-Therefore, our bounds can be used as a point of comparison for the optimality of a given public key, in the following sense.
-Say that, in order to forge a signature produced with a given multivariate digital signature scheme, one has to find a solution of a system of $m$ equations of degree $D$ in $n$ variables.
-Suppose that such a system is algebraically random.
-Our results provide an upper bound $B$ for the degree of regularity or the solving degree of such a system.
-Say that one can compute or estimate by a different method the degree of regularity or the solving degree of the specific system and suppose that this turns out to be $C$.
-Clearly, it must always be that $C\leq B$.
-However, how far $C$ is from $B$ gives us a measure of how close to optimal the digital signature scheme is, for the given choice of parameters.
-In fact, since our bounds on the degree of regularity are sharp, there are systems of $m$ equations of degree $D$ in $n$ variables whose degree of regularity is exactly $B$.
-In other words, if $B$ and $C$ are close, then there is not much space for improvement, since any system with the same parameters as our public key can have degree of regularity or solving degree at most $B$.
-If on the contrary $B$ and $C$ are far apart, then potentially there is a lot of space for finding a more robust system with the same parameters, since a system with those parameters can have degree of regularity or solving degree up to $B$.
+Polynomial systems play an important role within cryptography, as several cryptographic schemes rely on the difficulty of solving multivariate polynomial equations over finite fields, especially in the post‑quantum setting.
+Most notably, multivariate cryptography relies on the difficulty of solving multivariate polynomial equations over finite fields as its core security assumption.
+The public key of a multivariate system is a polynomial system, which is easy to evaluate, but is believed to be infeasible to solve without secret structural information, giving rise to candidate quantum‑resistant encryption and signature schemes.
+Beyond multivariate schemes, computational problems can often be modeled via polynomial systems, whose solutions provide a solution to the original problem.
+MinRank [15] is an example of such a problem, and the most efficient way of solving it currently consists of modeling it via a polynomial system [5, 2, 4].
+In algebraic cryptanalysis, block ciphers, hash functions, and other primitives are sometimes translated into large systems of polynomial equations, whose solutions allow an attacker to recover keys or find collisions, see e.g.
+[33, 40, 1, 22, 24, 34, 18].
+The security of many cryptographic schemes, therefore, relies on the complexity of solving the related polynomial system.
+The degree of regularity of a polynomial system provides an upper bound on the complexity of solving the system using Gröbner bases [38], and hence an upper bound on the security of the corresponding scheme.
+However, often the degree of regularity is hard to compute or even estimate.
+In practice, one cannot compute it directly for systems of cryptographic size, as this would require a Gröbner basis computation for values of the parameters for which it is considered infeasible.
+Depending on the system in question, the degree of regularity and other related invariants may also be hard to estimate theoretically.
+Due to these difficulties, cryptographers often rely on heuristic estimates.
+A widespread approach consists of assuming that a system is semiregular (see e.g.
+[19]), which provides an explicit formula for its degree of regularity [7].
+In this paper, we consider systems for which the degree of regularity is finite.
+These are exactly the systems for which the approach via the degree of regularity provides a non-trivial complexity bound.
+For polynomials in $n$ variables, the degree of regularity is finite if and only if the ideal generated by the top-degree part of the system contains a regular sequence of $n$ polynomials.
+For a system of polynomials of the same degree $D$, the degree of regularity is finite if and only if the regular sequence may be found in degree $D$.
+For such systems, in Theorem 23 we prove a sharp upper bound for the degree of regularity in terms of the number of equations, the number of variables, and the degree of the equations.
+In Theorem 33 we prove a similar result for a system of equations to which we add the field equations.
+Our bounds on the degree of regularity readily produce bounds on the complexity of computing a Gröbner basis, hence of solving, many systems that are of interest in cryptography.
+Bounds on the complexity produced in this way have the advantage of being widely applicable and the disadvantage of not always being close to the actual complexity of solving the systems to which they apply.
+In particular, the complexity estimates they produce are less sharp than those produced using the heuristic assumption that the system is semiregular.
+However, our bounds are proven and do not rely on any heuristic assumption.
+In particular, they apply to the situations where no semiregular sequence exists.
+This is the case for certain choices of the parameters and over a given field.
+On the other hand, since they are sharp, our bounds give us an idea of what security one can hope to achieve for a system with given parameters.
 The paper is structured as follows.
 Section 2 contains some useful preliminaries.
-In Section 3 we discuss the concept of randomness and propose two definitions of randomness for a polynomial system, in Definition 7 and Definition 12.
-Throughout the paper, we refer to our notion of randomness as algebraic randomness.
-We briefly review cryptographic semiregular sequences and semiregular sequences and show that the latter are algebraically random.
-In Section 4, after recalling some results from commutative algebra, we prove some bounds on the degree of regularity and the solving degree of an algebraically random system consisting of polynomials the same degree.
-The bounds appear in Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32.
-Finally, in Section 5 we show how to apply our results to check that the systems associated to GeMSS and Rainbow are not algebraically random.
+In Section 3 we briefly review cryptographic semiregular sequences and semiregular sequences, and we introduce the family of polynomial systems that are regular in a given degree $D$.
+These are the systems that we will focus on in this paper.
+In Section 4, after recalling some results from commutative algebra, we establish bounds on the degree of regularity and the solving degree of a polynomial system that is regular in degree $D$.
+If the field size is smaller than the degree $D$ of the equations, then we establish bounds on the degree of regularity and the solving degree of the polynomial system to which we add the field equations.
+The bounds appear in Theorem 23, Corollary 27, Proposition 31, and Theorem 33.
 2 Preliminaries
-Let $\mathbb{F}_{q}$ be the finite field of cardinality $q$ and $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the polynomial ring over $\mathbb{F}_{q}$ in $n$ variables.
-We start by recalling how a typical multivariate one-way function is constructed.
-Let $f_{1},\ldots,f_{m}\in R$, and consider the evaluation map
+Let $\mathbb{F}$ be an arbitrary field and let $\mathbb{F}_{q}$ be the finite field of cardinality $q$.
+Let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ be the ring of polynomials in $n$ variables with coefficients in the field $\mathbb{F}$.
+Even though our main motivation comes from cryptography, most of the results in this paper do not depend on the ground field and will be stated and proven over an arbitrary field.
+We start by recalling how a typical multivariate one-way function is constructed. Let $\mathbb{F}=\mathbb{F}_{q}$, $f_{1},\ldots,f_{m}\in R$, and consider the evaluation map
 $\begin{array}[]{cccc}\mathcal{F}:&\mathbb{F}_{q}^{n}&\rightarrow&\mathbb{F}_{q}^{m}\\ &\alpha=(\alpha_{1},\ldots,\alpha_{n})&\mapsto&(f_{1}(\alpha_{1},\ldots,\alpha_{n}),\ldots,f_{m}(\alpha_{1},\ldots,\alpha_{n}))\end{array}$
-To hide the structure of $\mathcal{F}$, we compose it with two random invertible linear maps $S:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ and $T:\mathbb{F}_{q}^{m}\rightarrow\mathbb{F}_{q}^{m}$.
-We obtain $\mathcal{P}=T\circ\mathcal{F}\circ S$, a set of $m$ polynomials $p_{1},\ldots,p_{m}$ in $n$ variables over $\mathbb{F}_{q}$.
+To hide the structure of $\mathcal{F}$, one composes it with two random invertible linear maps $S:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ and $T:\mathbb{F}_{q}^{m}\rightarrow\mathbb{F}_{q}^{m}$.
+One obtains $\mathcal{P}=T\circ\mathcal{F}\circ S$, a set of $m$ polynomials $p_{1},\ldots,p_{m}$ in $n$ variables over $\mathbb{F}_{q}$.
 The public key of the multivariate scheme is $\mathcal{P}=(p_{1},\ldots,p_{m})$ and the private key is $\{\mathcal{F},S,T\}$.
@@
 Finding inverse images with respect to $\mathcal{P}$ corresponds to computing solutions of multivariate polynomial systems.
-It is well-known that this problem can be solved by computing a Gröbner basis of the system, we refer the interested reader to [13] for more detail.
+It is well-known that this problem can be solved by computing a Gröbner basis of the system $\mathcal{P}$, we refer the interested reader to [16] for more detail.
 The first algorithm for computing Gröbner bases appeared in the doctoral thesis of Buchberger [13].
-Modern algorithms for computing Gröbner bases are based on liner algebra and are more efficient than Buchberger’s.
+Modern algorithms for computing Gröbner bases are based on liner algebra and are often more efficient than Buchberger’s.
 Examples of linear-algebra-based algorithms are $F_{4}$ [28], $F_{5}$ [29], the XL Algorithm [23], MutantXL [14], and their variants.
-In all of these systems, one computes the reduced row echelon form of the Macaulay matrix associated to the polynomial equations in a given degree, for one or more degrees.
-We now describe the main object of linear-algebra-based algorithms [6, Section 1.2].
-Fix a term order on $R$ and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subseteq R$ be a system of homogeneous polynomials.
-The columns of the homogeneous Macaulay matrix $M_{d}$ of $\mathcal{F}$ are labelled by the monomials of $R_{d}$ and arranged in decreasing order.
-The rows of $M_{d}$ are labelled by polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$.
-The entry $(i,j)$ of $M_{d}$ is the coefficient of the monomial of column $j$ within the polynomial corresponding to the $i$-th row.
-Let now $f_{1},\dots,f_{m}$ be arbitrary (not necessarily homogeneous) polynomials.
-The columns of the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}$ are labelled by the monomials of $R$ of degree $\leq d$, arranged in decreasing order.
+All these algorithms compute the reduced row echelon form of the Macaulay matrix of the system in a given degree, for one or more degrees.
+We now define Macaulay matrices, see also [8, Section 1.2].
+Fix a term order on $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ and let $\mathcal{F}=\{f_{1},\dots,f_{m}\}\subseteq R$ be a system of homogeneous polynomials.
+The columns of the homogeneous Macaulay matrix $M_{d}$ of $\mathcal{F}$ are labeled by the monomials of degree $d$ in $R$ and arranged in decreasing order.
+The rows of $M_{d}$ are labeled by polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$.
+The entry of $M_{d}$ in position $(i,j)$ is the coefficient of the monomial of column $j$ in the polynomial corresponding to the $i$-th row.
+Let $f_{1},\dots,f_{m}$ be arbitrary polynomials (not necessarily homogeneous).
+The columns of the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}$ are labeled by the monomials of $R$ of degree $\leq d$, arranged in decreasing order.
 The rows of $M_{\leq d}$ correspond to polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})\leq d$.
 The entries of $M_{\leq d}$ are defined as in the homogeneous case.
-The rationale behind the use of homogeneous Macaulay matrices for homogeneous systems is that, for a homogeneous system, the Macaulay matrix $M_{\leq d}$ is a block matrix with blocks $M_{d},\ldots,M_{0}$ as blocks.
-One computes the reduced row echelon form of the Macaulay matrix, or of its homogeneous version, in one or more degrees.
-For large enough degree, this produces a reduced Gröbner basis with respect to the chosen order.
-Some algorithms, as e.g.
-MutantXL, use a variation called mutant strategy in the non-homogeneous case: If the reduction of the Macaulay matrix $M_{\leq d}$ produces new polynomials $g_{1},\ldots,g_{\ell}$ of degree strictly smaller than $d$, one appends to the reduction of $M_{\leq d}$ the polynomials $m_{i,j}g_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}g_{j})\leq d$, then computes the reduced row echelon form again.
-Throughout the paper, we refer to the algorithms that employ the mutant strategy as mutant algorithms and to the others as standard algorithms.
-For a description of the basic version of these algorithms, we refer to [13, Section 3.1] and [14, Section 1].
-The computational complexity of computing the reduced row echelon form of the Macaulay matrices $M_{\leq d}$ and $M_{d}$ depends on their size, and hence on the degree $d$.
-This motivates the next definition.
+The rationale behind the use of homogeneous Macaulay matrices for homogeneous systems is that, for a homogeneous system, the Macaulay matrix $M_{\leq d}$ is a block matrix with blocks $M_{d},\ldots,M_{0}$.
+In order to compute a Gröbner basis of the system $\mathcal{F}$, one computes the reduced row echelon form of the corresponding Macaulay matrix in one or more degrees.
+If the system is homogeneous, one uses homogeneous Macaulay matrices.
+For a sufficiently large degree, this produces a reduced Gröbner basis with respect to the chosen order.
+Some algorithms, such as MutantXL, use a variation called mutant strategy in the non-homogeneous case: If row reduction of the Macaulay matrix $M_{\leq d}$ produces new polynomials $g_{1},\ldots,g_{\ell}$ of degree strictly smaller than $d$, one appends to the reduction of $M_{\leq d}$ the polynomials $m_{i,j}g_{j}$, for every $j$ and every monomial $m_{i,j}\in R$ such that $\deg(m_{i,j}g_{j})\leq d$.
+Then one computes the reduced row echelon form again.
+Throughout the paper, we refer to algorithms that employ the mutant strategy as mutant algorithms and to the others as standard algorithms.
+For a description of the basic version of these algorithms, we refer to [16, Section 3.1] and [17, Section 1].
+The complexity of computing the reduced row echelon form of the Macaulay matrices $M_{\leq d}$ and $M_{d}$ depends on their size, hence on the degree $d$. This motivates the next definition.
 Definition 1.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ and let $\tau$ be a term order on $R$.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ and let $\tau$ be a term order on $R$.
 The solving degree of $\mathcal{F}$ with respect to $\tau$ is the least degree $d$ such that Gaussian elimination on the Macaulay matrix $M_{\leq d}$ produces a $\tau$-Gröbner basis of $\mathcal{F}$.
@@
 When $\tau$ is the degree reverse lexicographic order, we omit the subscript $\tau$.
-In general, the solving degree is not invariant under coordinate change.
-Moreover, it may depend on the algorithm used to perform the Gröbner basis computation.
+It is well-known that the solving degree is not invariant under coordinate change.
+In addition, the solving degree may depend on the algorithm used to perform the Gröbner basis computation.
 In particular, for mutant algorithms it may be smaller than for standard ones.
 Finally, the solving degree depends on the choice of a term order on $R$.
-The complexity of linear-algebra-based algorithms is dominated by the cost of computing a degree reverse lexicographic Gröbner basis of the system, see [13, Sections 2 and 3] for more detail.
+The complexity of linear-algebra-based algorithms is dominated by the cost of computing a degree reverse lexicographic Gröbner basis of the system, see [16, Sections 2 and 3] for more detail.
 Therefore, an upper bound on the solving degree with respect to the reverse lexicographic order yields an upper bound on the complexity of computing a lexicographic Gröbner basis, hence on the complexity of solving the polynomial system.
 Let $I$ be a homogeneous ideal of $R$.
-For an integer $d\geq 0$, we denote by $I_{d}=I\cap R_{d}$ the $\mathbb{F}$-vector space of homogeneous polynomials of degree $d$ in $I$.
+For an integer $d\geq 0$, we denote by $R_{d}$ the vector space generated by the degree $d$ monomials of $R$.
+We denote by $I_{d}=I\cap R_{d}$ the $\mathbb{F}$-vector space of homogeneous polynomials of degree $d$ in $I$.
 For $g\in R$ a polynomial, we denote by $g^{\operatorname{\mathrm{top}}}$ the homogeneous part of $g$ of largest degree.
-E.g., if $g=x^{3}+2xy^{2}-y+1\in\mathbb{F}[x,y]$, then $g^{\operatorname{\mathrm{top}}}=x^{3}+2xy^{2}$.
-For a polynomial system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$, we denote by $\mathcal{F}^{\operatorname{\mathrm{top}}}\subseteq R$ the homogeneous system $\{f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\}$.
-Up to doing Gaussian elimination in a matrix whose rows correspond to $f_{1},\ldots,f_{m}$, we may suppose that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent.
-We will assume this throughout the paper.
+For example, if $g=x^{3}+2xy^{2}-y+1\in\mathbb{F}[x,y]$, then $g^{\operatorname{\mathrm{top}}}=x^{3}+2xy^{2}$.
+For a polynomial system $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$, we denote by $\mathcal{F}^{\operatorname{\mathrm{top}}}\subseteq R$ the homogeneous system $\{f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\}$.
+Up to performing Gaussian elimination on a matrix whose rows correspond to $f_{1},\ldots,f_{m}$, we may suppose that
+$f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent.
+Throughout the paper, we always make this assumption.
 The degree of regularity was introduced in [3, 7].
@@
 $d_{\operatorname{reg}}(\mathcal{F})=\left\{\begin{array}[]{cc}\min\{d\geq 0\mid(\mathcal{F}^{\mathrm{top}})_{d}=R_{d}\}&\mbox{ if }\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}\mbox{ for $d\gg 0$}\\ +\infty&\mbox{ otherwise.}\end{array}\right$.
-In the cryptographic literature, the degree of regularity is often used as a proxy for the solving degree.
-This is the case, e.g., in the specification documents of GeMSS [15].
-However, this does not always produce reliable estimates.
-In fact, there are examples in which the gap between the degree of regularity and the solving degree is large, see e.g.
-[10, Examples 3.2 and 3.3].
-A recent result by Semaev and Tenti [27, 28] however shows that, under suitable assumptions, the solving degree of a system with respect to a standard algorithm is at most twice the degree of regularity.
-Thanks to this result, an upper bound for the degree of regularity yields a proven upper bound for the solving degree.
-Theorem 3 ([28, Corollary 3.67] and [27, Theorem 2.1]).
-Let $\mathbb{F}=\mathbb{F}_{q}$.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m},x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\subseteq R$ be a polynomial system.
-If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then for standard algorithms
-$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2d_{\operatorname{reg}}(\mathcal{F})-2.$
-Notice that, in almost all systems of cryptographic interest, the field size and the degrees of the polynomials are relatively small. Therefore, one expects that Theorem 3 applies to such systems.
+In the cryptographic literature, the degree of regularity is often assumed to be finite and is used as a proxy for the solving degree.
+This is the case, e.g., in the GeMSS specification documents [19].
+However, this does not always produce reliable estimates, even in the case when the degree of regularity is finite.
+In fact, there are examples in which the degree of regularity is much smaller than the solving degree, see e.g.
+[12, Examples 3.2 and 3.3].
+However, a result by Semaev and Tenti [39, 41] shows that, under suitable assumptions, the solving degree of a system which contains the field equations is at most twice the degree of regularity.
+Due to this result, an upper bound for the degree of regularity yields a proven upper bound for the solving degree of standard algorithms.
+Theorem 3 ([41, Corollary 3.67] and [39, Theorem 2.1]).
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m},x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\subseteq R$ be a polynomial system.
+If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then
+$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F})\leq 2d_{\operatorname{reg}}(\mathcal{F})-2.$
+Notice that in most systems of cryptographic interest, the field size and the degrees of the polynomials are relatively small. Therefore, one expects that Theorem 3 applies to such systems.
 Another recent result by Salizzoni [38] shows that the solving degree of a mutant algorithm is at most the degree of regularity plus one, unless the system contains polynomials of large degree.
-Theorem 4 ([26, Proposition 3.10]).
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system, then for mutant algorithms
-$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{d_{\operatorname{reg}}(\mathcal{F})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}.$
+Theorem 4 ([38, Proposition 3.10]).
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system, then
+$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F})\leq\max\{d_{\operatorname{reg}}(\mathcal{F})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}.$
+Notice that, for both standard and mutant algorithms, it makes sense to assume that
+$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\geq\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}.$ (1)
+In fact, any equation of $\mathcal{F}$ of degree greater than the solving degree may be removed from the system without altering the Gröbner basis computation.
+This is of course delicate in practice, as one may not have an a priori estimate on the solving degree that allows them to eliminate equations from the system.
+Notice moreover that most multivariate systems associated to cryptographic schemes have low degree, so (1) is usually satisfied.
 Another algebraic invariant connected to the solving degree of a polynomial system is the Castelnuovo-Mumford regularity.
-We refer the reader to [13, Section 3.4] and [14] for its definition and a discussion on its relation with the other invariants.
-Remark 5.
-If $\mathbb{F}=\mathbb{F}_{q}$ and the system $\mathcal{F}$ contains the field equations $x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}$, then $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, therefore the degree of regularity is finite.
-Remark 6.
-If $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, then
+Throughout the paper, we denote by $\operatorname{reg}(I)$ the Castelnuovo-Mumford regularity of the ideal $I$.
+We refer the reader to [16, Section 3.4] and [17] for its definition and a discussion of its relationship with the other invariants.
+Notice that, if $d_{\operatorname{reg}}(\mathcal{F})<+\infty$, then
 $d_{\operatorname{reg}}(\mathcal{F})=\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}}).$
-Therefore:
-  • By Theorem 4
-    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}$
-    for mutant algorithms.
-  • If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then by Theorem 3
-    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})-2$
-    for standard algorithms.
-3 Random polynomial systems
-Consider a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables.
-Over a finite field, one often defines a random system as a polynomial system whose coefficients are chosen uniformly at random in the given field.
-In this paper, we propose a different definition which still captures the intuitive idea of randomness, while allowing us to estimate the degree of regularity of a random system.
-Over an infinite field, one may use the concept of genericity from algebraic geometry to define randomness.
-More precisely, fix a nonempty Zarisky-open subset of $\mathbb{P}^{\binom{n+d_{1}}{n}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}}{n}-1}$, where $\mathbb{P}^{t}$ denotes $t$-dimensional projective space.
-Define a random system as an element of that open set.
-This makes sense, since every nonempty Zarisky-open set is dense, hence a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables whose coefficients are chosen uniformly at random is generic with high probability according to this definition.
-The problem of extending this definition to a finite field is that, over a finite field, a nonempty Zarisky-open set is no longer dense, so the connection with our intuitive idea of randomness is lost.
-Nevertheless, for a finite field $\mathbb{F}_{q}$, we may define randomness using a Zariski-dense open set defined over the algebraic closure $\overline{\mathbb{F}_{q}}$.
-While it is not necessarily the case that almost every polynomial system of given degrees with coefficients in $\mathbb{F}_{q}$ is random, this is the case whenever $q$ is large enough, or if we consider a finite extension of $\mathbb{F}_{q}$ of large enough cardinality.
+3 Systems that contain a regular sequence in a given degree
+Consider a system $\mathcal{F}$ of $m$ equations in $n$ variables.
+As we discussed in the previous section, whenever the degree of regularity of $\mathcal{F}$ is finite, an upper bound on it directly translates into an upper bound on the solving degree of the system, hence on the complexity of computing its solutions.
+In this paper, we prove upper bounds on the degree of regularity of a system $\mathcal{F}$ such that the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}$ contains a regular sequence of polynomials of degree $D$.
+In the rest of this section, we discuss the property of containing a regular sequence in a given degree and in which sense this is a generic property.
+The next definition formalizes the correspondence between homogeneous systems and points in a product of projective spaces.
+Definition 5.
+Let $\mathbb{F}$ be a field, and let $\overline{\mathbb{F}}$ be its algebraic closure.
+Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$.
+Let $d_{1},\ldots,d_{m}$ be positive integers.
+For a homogeneous polynomial $f$, denote by $f]\in\mathbb{P}^{\binom{n+\deg(f)-1}{n-1}-1}$ the projective point whose coordinates are the coefficients of $f$.
+A homogeneous system $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$ for $1\leq i\leq m$ corresponds to a point $([f_{1}],\ldots,[f_{m}])\in\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$.
+A random system of degrees $d_{1},\ldots,d_{m}$ over a field $\mathbb{F}$ is a system of polynomials of the given degrees whose coefficients are chosen uniformly at random in the given field.
+Similarly, a random homogeneous system of degrees $d_{1},\ldots,d_{m}$ is a system of homogeneous polynomials of the given degrees, whose coefficients are chosen uniformly at random.
+The concept of random system is used extensively in the cryptographic literature, where random instances of structured polynomial systems are often assumed to behave as random systems.
+Over an infinite field, a generic property in the sense of algebraic geometry is by definition a property which holds over a nonempty Zarisky-open set.
+We refer to [16, Section 1.3] for a short reminder on the Zarisky topology.
+Since every nonempty Zarisky-open set is dense, the property holds for “almost every system”.
+Therefore, over an infinite field, genericity is the algebraic-geometric analog of being true with high probability.
+Over a finite field $\mathbb{F}_{q}$, however, estimating the probability that a system has a given property amounts to estimating the number of $\mathbb{F}_{q}$-rational points of the open set that corresponds to the property over the algebraic closure, a task which is often far from trivial.
+Nevertheless, if $q$ is large enough (or if we consider a finite extension of $\mathbb{F}_{q}$ of large enough cardinality), a random homogeneous system belongs to any given nonempty Zarisky-open set with high probability.
+For the convenience of the reader, we show how to deduce this from known results.
+For simplicity, we only state the argument for one projective space, but the same argument can be applied to a product of projective spaces.
+Proposition 6.
+Let $\mathbb{P}^{t}$ be the $t$-dimensional projective space over $\bar{\mathbb{F}_{q}}$.
+Let $\mathcal{U}\subseteq\mathbb{P}^{t}$ be a nonempty Zarisky-open set defined over $\mathbb{F}_{q}$.
+If $q$ is sufficiently large, then a randomly chosen $\mathbb{F}_{q}$-rational point $P\in\mathbb{P}^{t}$ belongs to $\mathcal{U}$ with high probability.
+Proof.
+Let $\mathcal{X}$ be the set-theoretic complement of $\mathcal{U}$ in $\mathbb{P}^{t}$.
+Let $Q$ be a power of $q$.
+It suffices to show that the probability that a randomly chosen $\mathbb{F}_{Q}$-rational point belongs to $\mathcal{X}$ tends to zero as $Q$ tends to infinity.
+Since $\mathcal{X}$ is a projective algebraic variety, it follows from [25, Theorem 3.1] that the number $|\mathcal{X}(\mathbb{F}_{Q})|$ of $\mathbb{F}_{Q}$-rational points of $\mathcal{X}$ is bounded from above by a polynomial $p(Q)$, whose coefficients depend on the degrees and dimensions of the irreducible components of $\mathcal{X}$.
+In particular,
+$\deg p(Q)\leq\max\{\dim(\mathcal{X}),2\dim(\mathcal{X})-t\}\leq t-1.$
+Since the number of $\mathbb{F}_{Q}$-rational points of $\mathbb{P}^{t}$ is $\sum_{i=0}^{t}Q^{i}$, then the probability that an $\mathbb{F}_{Q}$-rational belongs to $\mathcal{X}$ is bounded from above by
+[…]
+Hence, the probability tends to zero as $Q$ tends to infinity. ∎
+In the cryptographic literature, random polynomial systems are often assumed to be cryptographic semiregular sequences, see e.g.
+[3, 6].
+This is the case in the cryptoanalysis of several systems, as e.g.
+[19].
+The next definition appears in [3, Definition 3.2.1 and Definition 3.2.4], [7, Definition 5], and [9, Definition 5 and Definition 9].
+Definition 7.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a homogeneous system.
+If $\mathbb{F}\neq\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F})$, we have $g_{i}\in(f_{1},\ldots,f_{i-1})$.
+If $\mathbb{F}=\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2},\ldots,x_{n}^{2}\})$, we have $g_{i}\in(f_{1},\ldots,f_{i})$.
+A sequence of polynomials (not necessarily homogeneous) $f_{1},\ldots,f_{m}$ is a cryptographic semiregular sequence if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are a cryptographic semiregular sequence.
+In this paper, we use the expression cryptographic semiregular sequence in order to distinguish the concept of semiregularity used in the cryptographic literature from the concept of semiregularity originally introduced by Pardue [35, 36], which inspired it.
+The original definition by Pardue is given over an infinite field $\mathbb{F}$.
+As we are interested in dealing with finite fields, we extend it to an arbitrary field in the natural way.
+Definition 8.
+Let $\mathbb{F}$ be a field, and let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$.
+Let $I$ be a homogeneous ideal, and let $A=R/I$.
+A polynomial $f\in R_{d}$ is semiregular on $A$ if for every $e\geq d$, the vector space map $A_{e-d}\rightarrow A_{e}$ given by multiplication by $f$ has maximal rank (that is, it is either injective or surjective).
+A sequence of homogeneous polynomials $f_{1},\ldots,f_{m}$ is a semiregular sequence if $f_{i}$ is semiregular on $A/(f_{1},\ldots,f_{i-1})$ for all $1\leq i\leq m$.
+It follows from [36, Proposition 1] that, if $\mathbb{F}\neq\mathbb{F}_{2}$, then a semiregular sequence is also a cryptographic semiregular sequence.
+The converse does not hold, as shown in [36], see the example just below [36, Proposition 1].
+The next proposition presents a simple situation in which cryptographic semiregular sequences and semiregular sequences coincide.
+Proposition 9.
+Let $q>2$ and let $f_{1},\ldots,f_{n+1}\in R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be polynomials of degree $d_{1},\ldots,d_{n+1}$, respectively.
+Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}}$ is a regular sequence.
+The sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular if and only if the sequence $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular.
+In particular, the sequence $x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n},f$ is cryptographic semiregular if and only if the sequence $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular.
+Proof.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{n+1}\}$.
+It suffices to show that if the sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then the sequence $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular.
+Since $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}}$ is a regular sequence, it suffices to show that $f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}})$.
+If $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then by [9, Proposition 6] the Hilbert series of $R/(\mathcal{F}^{\operatorname{\mathrm{top}}})$ is $(1-z^{d_{n+1}})\prod_{i=1}^{n}(1+z+\ldots+z^{d_{i}-1})]$.
+Here, for $p(z)=\sum_{i=0}^{\infty}p_{i}z^{i}\in\mathbb{Z}[[z]]$, we denote by $\delta(p)=\min\{i\geq 0\mid p_{i}\leq 0\}-1$ and define $p(z)]=\sum_{i=0}^{\delta(p)}p_{i}z^{i}$.
+Then $f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}})$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a semiregular sequence by [36, Proposition 1].
+∎
+Pardue in [35, 36] shows that Fröberg’s Conjecture [30], a conjecture which has attracted a lot of attention within the commutative algebra community and is widely believed to hold, is equivalent to the following
+Conjecture 10.
+Let $\mathbb{F}$ be an infinite field. A generic sequence of polynomials of degrees $d_{1},\ldots,d_{m}$ in $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ is semiregular.
+In other words, Fröberg’s Conjecture is equivalent to the statement that the set of semiregular sequences of polynomials of given degrees contains a dense Zarisky-open set.
+If the conjecture is true, then a sequence of polynomials of given degrees is semiregular with high probability, provided that the ground field has large enough cardinality.
+It follows that, if $\mathbb{F}=\mathbb{F}_{q}$ with $q\gg 0$ and Fröberg’s Conjecture holds, then a sequence of polynomials of given degrees is a cryptographic semiregular sequence with high probability.
+In addition, for $q\gg 0$ most cryptographic semiregular sequences are also semiregular sequences, as the set of semiregular sequences conjecturally contains a dense open set.
+In [3, Section 3.2], Bardet conjectures that a sequence of polynomials with coefficients in $\mathbb{F}_{2}$ is cryptographic semiregular with high probability.
+This conjecture is motivated by experimental evidence, see also [6, Conjecture 2].
+The conjecture was later disproved by Hodges, Molina, and Schlather, who in [31] prove that there are choices of the parameters for which no cryptographic semiregular sequence exists over $\mathbb{F}_{2}$.
+This is the case, e.g., for $m=1$ and $n>3d_{1}$.
+In this paper, we study the Zariski-open set consisting of systems of $m$ polynomials in $n$ variables which generate an ideal that contains a regular sequence of $n$ polynomials in a given degree $D$.
 Definition 11.
@@
 Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$.
-Let $d_{1},\ldots,d_{m}$ be positive integers and let $\mathcal{U}\subseteq\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ be a nonempty Zariski-open set.
-For a homogeneous polynomial $f$, denote by $[f]\in\mathbb{P}^{\binom{n+\deg(f)-1}{n-1}-1}$ the projective point, whose coordinates are the coefficients of $f$.
-A homogeneous system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$ for $1\leq i\leq m$ is generic with respect to $\mathcal{U}$ if $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$.
-It is generic if it is generic with respect to $\mathcal{U}$, for some nonempty Zariski-open set $\mathcal{U}$.
-An arbitrary system $\mathcal{F}$ is generic or generic with respect to $\mathcal{U}$ if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is.
-In the cryptographic literature, random sequences of polynomials are often assumed to be cryptographic semiregular sequences, see e.g.
-[2, 4].
-This is the case in the cryptoanalysis of several systems, as e.g.
-[15].
-The next definition appears in [2, Definition 3.2.1 and Definition 3.2.4][5, Definition 5], and [7, Definition 5 and Definition 9].
-Definition 8.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a homogeneous system.
-If $\mathbb{F}\neq\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F})$, one has that $g_{i}\in(f_{1},\ldots,f_{i-1})$.
-If $\mathbb{F}=\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2},\ldots,x_{n}^{2}\})$, one has that $g_{i}\in(f_{1},\ldots,f_{i})$.
-Arbitrary polynomials $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are a cryptographic semiregular sequence.
-In this paper, we use the word cryptographic semiregular sequence in order to distinguish the concept of semiregularity used in the cryptographic literature from the concept of semiregularity originally introduced by Pardue [24, 25], which inspired it.
-The original definition by Pardue is given over an infinite field $\mathbb{F}$.
-As we are interested also in dealing with finite fields, we extend it in the natural way.
-Definition 9.
-Let $\mathbb{F}$ be an infinite field and let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$.
-Let $I$ be a homogeneous ideal and let $A=R/I$.
-A polynomial $f\in R_{d}$ is semiregular on $A$ if for every $e\geq d$, the vector space map $A_{e-d}\rightarrow A_{e}$ given by multiplication by $f$ has maximal rank (that is, it is either injective or surjective).
-If $\mathbb{F}$ is a finite field, let $\overline{R}=\overline{\mathbb{F}}[x_{1},\ldots,x_{n}]$.
-Then $f$ is semiregular on $A$ if it is semiregular on $\overline{R}/I\overline{R}$.
-A sequence of homogeneous polynomials $f_{1},\ldots,f_{m}$ is a semiregular sequence if $f_{i}$ is semiregular on $A/(f_{1},\ldots,f_{i-1})$ for all $1\leq i\leq m$.
-It follows from [25, Proposition 1] that, if $\mathbb{F}\neq\mathbb{F}_{2}$, then a semiregular sequence is also a cryptographic semiregular sequence.
-The converse does not hold, as shown in [25], see the example just below [25, Proposition 1].
-The next proposition presents a simple situation in which cryptographic semiregular sequences and semiregular sequences coincide.
+Fix $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$.
+Let $\mathcal{U}_{D}\subseteq\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ be the set of $([g_{1}],\ldots,[g_{m}])\in\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ such that the ideal $(g_{1},\ldots,g_{m})\subseteq\overline{\mathbb{F}}[x_{1},\ldots,x_{n}]$ contains a regular sequence of $n$ polynomials of degree $D$.
+We say that the system $\mathcal{F}$ is regular in degree $D$ if $([f_{1}^{\operatorname{\mathrm{top}}}],\ldots,[f_{m}^{\operatorname{\mathrm{top}}}])\in\mathcal{U}_{D}$.
+Fix $m\geq n\geq 1$, $1\leq d_{1}\leq\ldots\leq d_{m}$, and $\mathbb{F}$ a field.
+If $D<d_{n}$, then $\mathcal{U}_{D}=\emptyset$.
+In fact, let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$.
+Then $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})_{D}=(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n-1}^{\operatorname{\mathrm{top}}})_{D}$ cannot contain a regular sequence of $n$ polynomials.
+Therefore, the system $\mathcal{F}$ can be regular in degree $D$, only if $D\geq d_{n}$.
+The next proposition provides some useful facts connected to the property of being regular in a given degree.
+The statement follows from standard facts from commutative algebra.
+Recall that a system defined over a field $\mathbb{F}$ is zero-dimensional if it has finitely many solutions over $\overline{\mathbb{F}}$.
 Proposition 12.
-Let $q>2$ and let $f_{1}=x_{1}^{q}-x_{1},\ldots,f_{n}=x_{n}^{q}-x_{n},f_{n+1}=f\in R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$.
-The sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular if and only if the sequence $f_{1}^{\operatorname{\mathrm{top}}}=x_{1}^{q},\ldots,$ $f_{n}^{\operatorname{\mathrm{top}}}=x_{n}^{q},f_{n+1}^{\operatorname{\mathrm{top}}}=f^{\operatorname{\mathrm{top}}}$ is semiregular.
-Proof.
-It suffices to show that, if the sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then the sequence $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular.
-We start by observing that the sequence $x_{1}^{q},\ldots,x_{n}^{q}$ is regular.
-Therefore, in order to show that $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular, it suffices to show that $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$.
-If $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular and $d=\deg(f)$, then the Hilbert series of $R/(\mathcal{F}^{\operatorname{\mathrm{top}}})$ is $[(1-z^{d})(1+z+\ldots+z^{q-1})^{n}]$ by [7, Proposition 6].
-Here for $p(z)=\sum_{i=0}^{\infty}p_{i}z^{i}\in\mathbb{Z}[[z]]$, we denote by $\delta(p)=\min\{i\geq 0\mid p_{i}\leq 0\}-1$ and define $[p(z)]=\sum_{i=0}^{\delta(p)}p_{i}z^{i}$.
-Then $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a semiregular sequence by [25, Proposition 1].
-∎
-Pardue in [24, 25] shows that Fröberg’s Conjecture [22], a conjecture which has attracted a lot of attention within the commutative algebra community and that is widely believed to hold, is equivalent to the following
-Conjecture 11.
-Let $\mathbb{F}$ be an infinite field. A generic sequence of polynomials of degrees $d_{1},\ldots,d_{m}$ in $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ is semiregular.
-In other words, Fröberg conjectures that the set of semiregular sequences of polynomials of given degrees contains a dense Zarisky-open set.
-If the conjecture is true, then a sequence of polynomials of given degrees is semiregular with high probability, provided that the ground field has large enough cardinality.
-It follows that, if $\mathbb{F}=\mathbb{F}_{q}$ with $q\gg 0$ and Fröberg’s Conjecture holds, then a sequence of polynomials of given degrees is a cryptographic semiregular sequence with high probability.
-In addition, most cryptographic semiregular sequences are also semiregular sequences, as the set of semiregular sequences conjecturally contains a dense open set.
-In [2, Section 3.2], Bardet conjectures that a sequence of polynomials with coefficients in $\mathbb{F}_{2}$ is cryptographic semiregular with high probability.
-This conjecture is motivated by experimental evidence, see also [4, Conjecture 2].
-The conjecture was later disproved by Hodges, Molina, and Schlather, who in [23] prove that there are choices of the parameters for which no cryptographic semiregular sequence exists over $\mathbb{F}_{2}$.
-This is the case, e.g., for $m=1$ and $n>3d_{1}$.
-In the sequel, we propose a notion of randomness which applies to any choice of the system parameters.
-In this paper, we propose two dense Zariski-open sets, which can be used to formalize the intuitive idea of a random system.
-The set $\mathcal{V}$ corresponds to systems of $m$ polynomials in $n$ variables which contain a regular sequence of length $n$, while the set $\mathcal{U}$ parametrizes systems of $m$ polynomials which contain a regular sequence of $n$ polynomials of the smallest possible degrees.
-Notice that $\mathcal{V}$ contains $\mathcal{U}$ and, if $m\geq n$, it also contains the set of cryptographic semiregular sequences and that of semiregular sequences.
-Definition 12.
-Fix $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$.
-For any multiset $\Delta$ of cardinality $n$ contained in the multiset $\{d_{1},\ldots,d_{m}\}$, let $\mathcal{U}_{\Delta}$ be the subset of $\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ whose points correspond to polynomials $f_{1},\ldots,f_{m}\in R$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ contains a regular sequence in the degrees of $\Delta$.
-Let
-$\mathcal{U}=\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}\;\;\;\mbox{ and }\;\;\;\mathcal{V}=\bigcup_{\Delta\subseteq\{d_{1},\ldots,d_{m}\},|\Delta|=n}\mathcal{U}_{\Delta}.$
-An algebraically random system of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n$ variables is a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ such that $\deg(f_{i})=d_{i}$ for all $1\leq i\leq m$ and $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$.
-It is well-known that $\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}$ is a dense open set for any choice of $1\leq d_{1}\leq\ldots\leq d_{n}$, if $m=n$.
-This implies that any $\mathcal{U}_{\Delta}$ as in Definition 12 is a dense open set, hence also $\mathcal{U}$ and $\mathcal{V}$ are.
-Notice that the degree of regularity is finite for all systems in $\mathcal{U}_{\Delta}$ and for any choice of $\Delta$, therefore it is also finite for the systems in $\mathcal{U}$ and $\mathcal{V}$.
-In particular, the degree of regularity of an algebraically random system is finite.
-Notice moreover that, if $\mathcal{F}$ is homogeneous, then $\mathcal{V}$ is the set of polynomial systems of degrees $d_{1},\ldots,d_{m}$ for which the degree of regularity is finite.
-Notice moreover that, if $m=n$, an algebraically random system is a system for which $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence.
-Since this case is well-studied in the cryptographic literature, in the sequel we often assume $m>n$.
+Let $\mathbb{F}$ be a field, and let $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$. The following are equivalent:
+  1. 1.
+    $d_{\operatorname{reg}}(\mathcal{F})<+\infty$,
+  2. 2.
+    $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is zero-dimensional,
+  3. 3.
+    $\mathcal{F}^{\operatorname{\mathrm{top}}}\in\mathcal{U}_{D}$ for some $D\geq 0$,
+  4. 4.
+    $\mathcal{F}^{\operatorname{\mathrm{top}}}\in\mathcal{U}_{D}$ for all $D\geq d_{m}$.
+In particular, if $d_{1}=\ldots=d_{m}=D$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is zero-dimensional, then $\mathcal{F}$ is regular in degree $D$.
+It follows from Proposition 12 that, if $m=n$, a system $\mathcal{F}$ has finite degree of regularity if and only if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence.
+Since this case is well-studied, in the sequel we often focus on the case where $m>n$.
+Notice moreover that, if $\mathcal{F}$ is cryptographic semiregular, then it satisfies the equivalent conditions of Proposition 12.
+In other words, every cryptographic semiregular system is regular in degree $D$ for every $D\geq d_{m}$ (and possibly in smaller degree).
+However, the converse does not hold.
+The statement that every cryptographic semiregular sequence is regular in some degree also implies that being regular in degree $D$ is a generic property, conditionally to Conjecture 10.
+More precisely, Proposition 12 implies that, if $D\geq d_{n}$, then $\mathcal{U}_{D}$ contains a dense open set of $\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$.
+Even over a finite field, one advantage of considering systems which are regular in degree $D$ is that, unlike cryptographic semiregular systems [31], they exist over every field and for any choice of $m\geq n$ and $D\geq d_{n}$.
+In other words, if the set $\mathcal{U}_{D}$ is nonempty over $\overline{\mathbb{F}_{q}}$, then it always contains $\mathbb{F}_{q}$-rational points.
+Example 13.
+The set $\mathcal{U}_{D}$ contains $\mathbb{F}_{q}$-rational points for any choice of $n,m,q$ and $1\leq d_{1}\leq\ldots\leq d_{m}$ with $n\leq m$ and $D\geq d_{n}$.
+This follows from the existence of homogeneous regular sequences of any given degrees in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$.
+Some homogeneous regular sequences in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{n}$ are for example
+$x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},$
+where $g_{i}\in\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}]_{d_{i}}$. It follows that the system
+$\{x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},f_{n+1},\ldots,f_{m}\}$
+is regular in degree $D$ for every $D\geq d_{n}$ and for any $f_{n+1},\ldots,f_{m}$ of the appropriate degrees.
+We stress that Proposition 12 applies to most zero-dimensional systems, since for most zero-dimensional systems over an algebraically closed field, the top degree part of the system $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is also zero-dimensional.
+This statement can be formalized using the algebraic geometric language of dense open sets.
+Nevertheless, there exist zero-dimensional systems $\mathcal{F}$ such that $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is not zero-dimensional.
+Example 14.
+Let $f_{1}=x^{3}$, $f_{2}=xy^{2}+y^{2}$, $f_{3}=x^{2}y$.
+The system is zero-dimensional, as $\sqrt{(f_{1},f_{2},f_{3})}=(x,y)$.
+However, the system is not regular in degree $3$, since $(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})=(x^{3},x^{2}y,xy^{2})$ does not contain a regular sequence in degree $3$.
+In fact, $(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})$ is not even zero-dimensional, as $\sqrt{(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})}=(x)$.
+Example 15.
+The system $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ is regular in degree $q$, for any system $\mathcal{F}$.
+In fact, $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}},x_{1}^{q},\ldots,x_{n}^{q})_{q}$ contains the regular sequence $x_{1}^{q},\ldots,x_{n}^{q}$.
+4 Bounds on the degree of regularity
+Throughout the section, $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ denotes a system of $m$ polynomials in $n$ variables.
+We always assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent and $m\geq n$.
+Notice that
+$d_{\operatorname{reg}}(\mathcal{F})\geq d_{n}$
+by definition.
+In this section, we establish an upper bound for the degree of regularity of a system $\mathcal{F}$ of polynomials of equal degree $D$, under the assumption that $\mathcal{F}$ is regular in degree $D$.
+We also prove an upper bound for the degree of regularity of a system $\mathcal{F}$ of polynomials of equal degree $D$, for $D\geq q$.
+In combination with Theorem 3 and Theorem 4, these results provide us with upper bounds for the solving degree of $\mathcal{F}$.
 Remark 16.
-For any system $\mathcal{F}$ of equations of degree at least $q$, one has
-$\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}.$
-Unlike semiregular systems [23], algebraically random systems exist for any choice of the parameters and over every finite field.
-Remark 14.
-Algebraic random systems over $\mathbb{F}_{q}$ exist for any choice of $n,m$ and $1\leq d_{1}\leq\ldots\leq d_{m}$.
-This corresponds to the existence of regular sequences of any given degrees in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, as any system of $m$ equations in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ which contains a regular sequence in degrees $d_{1},\ldots,d_{n}$ is algebraically random.
-Some regular sequences in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{n}$ are for example
-$x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},$
-where $g_{i}\in\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}]$ and $\deg(g_{i})\leq d_{i}$.
-4 The degree of regularity of a random system
-In this section, we establish an upper bound for the degree of regularity of an algebraically random system $\mathcal{F}$ consisting of $m$ polynomials of equal degree $D$.
-In combination with Theorem 3 and Theorem 4, this provides us with an upper bound for the solving degree of a system of algebraically random polynomials of the same degree.
-Remark 15.
-Notice that $\mathcal{F}\in\mathcal{U}_{\{d,\ldots,d\}}$ for a given $d>0$ if and only if $T\circ\mathcal{F}\circ S\in\mathcal{U}_{\{d,\ldots,d\}}$.
-In other words, when deciding whether a system of polynomials of equal degree is algebraically random, one can safely ignore the random linear transformations $S$ and $T$ used to disguise the internal system $\mathcal{F}$.
-This also shows that, for system whose equations are all of the same degree, being an algebraically random system is an intrinsic property of the system and it is not affected by the invertible linear transformations used to disguise the system.
-In [10, Section 4], the authors provide an upper bound for the degree of regularity of a system of quadratic polynomials which contains a regular sequence.
+Let $\mathcal{F}$ be a system of polynomials with $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ linearly independent of degree $D$.
+Let $S$ and $T$ be invertible square matrices of size $n$ and $m$.
+Then $(T\circ\mathcal{F}\circ S)^{\operatorname{\mathrm{top}}}=T\circ\mathcal{F}^{\operatorname{\mathrm{top}}}\circ S$, so $\mathcal{F}$ is regular in degree $D$ if and only if $T\circ\mathcal{F}\circ S$ is regular in degree $D$.
+In other words, when deciding whether a system of polynomials of equal degree $D$ is regular in degree $D$, one can ignore the random linear transformations $S$ and $T$ used to disguise the internal system $\mathcal{F}$.
+In [12, Section 4], the authors provide an upper bound for the degree of regularity of a system of quadratic polynomials that contains a regular sequence.
 In this section, we follow the same basic approach and extend it to systems of polynomials of the same degree and systems of polynomials of the same degree to which one adds the field equations.
-The cases that we treat in this paper are technically more challenging and require the use of more sophisticated results from commutative algebra.
+The cases we treat in this paper are technically more challenging and require the use of more sophisticated results from commutative algebra.
 We start by introducing the family of lex-segment ideals.
-A conjecture by Eisenbud, Green, and Harris will allow us to reduce to these ideals, when estimating the regularity of ideals generated by algebraically random systems.
+A conjecture by Eisenbud, Green, and Harris will allow us to reduce to these ideals, when estimating the regularity of ideals generated by systems which are regular in degree $D$, for some $D>0$.
 Throughout the section, we fix the lexicographic term order on $R$ with $x_{1}>x_{2}>\ldots>x_{n}$.
@@
 An ideal $\mathcal{L}\subseteq R$ is a $(c_{1},\ldots,c_{n};C)$-LexPlusPowers (LPP) ideal if $\mathcal{L}=(x_{1}^{c_{1}},\ldots,x_{n}^{c_{n}})+L$, where $L$ is a lex-segment ideal generated in degree $C$.
-Notation 17.
+Notation 18.
 Let $I\subseteq R$ be a homogeneous ideal containing a regular sequence of polynomials of degrees $c_{1}\leq\ldots\leq c_{n}$.
@@
 The $(1,3,4;3)$-LPP ideal $\mathcal{L}$ such that $\dim(I_{3})=7=\dim(\mathcal{L}_{3})$ is $\mathcal{L}=(x,y^{3},z^{4})$.
-For any lex-segment ideal $L$ generated in degree $3$ with $\dim(L_{3})\leq 7$, one has $L\subseteq\mathcal{L}$. Hence one may write
+For any lex-segment ideal $L$ generated in degree $3$ with $\dim(L_{3})\leq 7$, one has $L\subseteq\mathcal{L}$. Hence, one may write
 $\mathcal{L}=(x,y^{3},z^{4})+L$
-as in Definition 16 and choose e.g. $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2})$, or $L=(x^{3})$. However, Notation 17 prescribes that we choose the largest $L$ with respect to containment, i.e.
+as in Definition 17 and choose, e.g., $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2})$, or $L=(x^{3})$. However, Notation 18 prescribes that we choose the largest $L$ with respect to containment, i.e.,
 $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2},y^{3})$
 is the lex-segment ideal generated by the first $7$ cubic monomials in the lexicographic order.
-The next conjecture appears as [19, Conjecture ($V_{m}$)].
-It has been settled in several cases and it is widely believed to hold within the commutative algebra community.
+The next conjecture appears as [27, Conjecture ($V_{m}$)].
+See also [37, Conjecture 2.6] for an equivalent formulation, that is closer to the formulation that we adopt here.
+The conjecture was settled in several cases and it is widely believed to hold within the commutative algebra community.
 For an introduction to the conjecture and an excellent survey of known cases, we refer the interested reader to [21].
@@
 for all $C\geq c_{n}$.
-In order to estimate the degree of regularity of our systems, we use the following result by Caviglia and De Stefani.
-Proposition 20 ([16, Lemma 2.3]).
+The conjecture is motivated by the following observation.
+A classical result by Macaulay shows that, among all homogeneous ideals with the same Hilbert function in a given degree $C$, the lex-segment generated in degree $C$ has the smallest Hilbert function in the next degree.
+If one restricts to ideals that contain a regular sequence in given degrees, the natural analog of the lex-segment is the LPP ideal from Notation 18.
+For ideals that contain a regular sequence of maximum cardinality, minimum growth of the Hilbert function corresponds to maximum Castelnuovo-Mumford regularity.
+In order to compute the degree of regularity of an LPP ideal, we use the following result by Caviglia and De Stefani.
+Proposition 21 ([20, Lemma 2.3]).
 Let $c_{1}\leq\ldots\leq c_{n}$ and $2\leq C\leq\sum_{i=1}^{n}(c_{i}-1)$.
@@
 $\operatorname{reg}(\mathcal{L})=t_{k}+\sum_{i=k+1}^{n}(c_{i}-1).$
-Our first result is an explicit bound for the degree of regularity of an algebraically random system of polynomials of the same degree.
+Our first result is an explicit upper bound for the degree of regularity of a system of polynomials that is regular in degree $D$.
 In order to make the proof more readable, we introduce the following notation.
-Notation 21.
+Notation 22.
 If $u\in R_{D}$ is a monomial that only involves the variables $x_{k},\ldots,x_{n}$ and has degree $a$ in $x_{k}$, we say that $u$ is a $(D,k,a)$-type monomial.
-In the next theorem, we provide an explicit formula for the degree of regularity of an LPP ideal with given parameters.
-Thanks to Conjecture 19, this yields an upper bound for the degree of regularity of an algebraically random system of polynomials of the same degree.
-In the statement of the theorem, $\sigma_{k,t}$ is the position of the smallest $(D,k,D-t)$-type monomial in the ordered list of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$, sorted in decreasing lexicographic order.
-In particular, $\sigma_{1,0}=0$ and $\sigma_{n-1,D-1}$ is the number of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$.
+In the next theorem, we provide an explicit upper bound for the degree of regularity of a system that is regular in degree $D$.
+The result is conditional on Conjecture 20.
+Notice that the result can be applied to practically any cryptographic system of equations of the same degree.
+In fact, even if the system of interest is not zero-dimensional, one can assign random values to some of the variables in order to obtain a zero-dimensional system, which is then solved for the remaining variables.
 Theorem 23.
 Assume that Conjecture 20 holds.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system and assume without loss of generality that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$.
-If $m=n$, then let $k=0$, $t=D-1$.
-Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
-$\sigma_{k,t}=\sum_{i=1}^{k}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}-\sum_{j=t+1}^{D-1}\binom{n-k-1+j}{j}.$
-If $\mathcal{F}$ is an algebraically random polynomial system, then
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system and assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$.
+If $m=n$, then let $k=0$, t=D1t=D-1.
+Else, let $1\leq k\leq n-1$ and $1\leq t\leq D-1$ be such that mnm-n belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
+$\sigma_{k,t}=\sum_{i=1}^{k}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}-\sum_{j=t+1}^{D-1}\binom{n-k-1+j}{j}\;\;\mbox{ for }\;0\leq t\leq D-1.$
+If $d_{\operatorname{reg}}(\mathcal{F})<+\infty$, then
 $d_{\operatorname{reg}}(\mathcal{F})\leq(D-t)+(n-k)(D-1).$
@@
 $\dim\left(\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle\right)=m-n.$
-For $11\leq k\leq n-$ 1 and $10\leq t\leq D-$ 1, the number of $(D,k,D-t)$-type monomials is
-$\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]_{t})=\binom{n-k-1+t}{t},$
+In order to apply Proposition 21, we need to determine the value of $k$ and $t_{k}$ for the monomial in position mnm-n in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$.
+As we will show, these values are the unique $k$ and $t_{k}$ such that mnm-n belongs to the interval $(\sigma_{k,D-t_{k}-1},\sigma_{k,D-t_{k}}]$.
+For $1\leq k\leq n-1$ and $0\leq t\leq D-1$, the number of (D,k,Dt)(D,k,D-t)-type monomials is
+$\dim(\mathbb{F}[x_{k+1},\ldots,x_{n}]_{t})=\binom{n-k-1+t}{t}$,
 hence
-$\begin{split}\sigma_{k,t}&=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}+\sum_{j=1}^{t}\binom{n-k-1+j}{j}\\ &=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\dim(\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}])_{j}+\sum_{j=1}^{t}\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}])_{j}\end{split}$
-is the number of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ different from $x_{1}^{D},\ldots,x_{n}^{D}$ and bigger than or equal to $x_{k}^{D-t}x_{n}^{t}$, the smallest $(D,k,D-t)$-type monomial.
-In other words, the monomial $x_{k}^{D-t}x_{n}^{t}$ is in position $\sigma_{k,t}$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$.
-Notice moreover that $\sigma_{1,0}=0$ and $\sigma_{k,0}=\sigma_{k-1,D-1}$ for $12\leq k\leq n-$ 1.
-If $u$ is the smallest monomial in $\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle$, then $u$ is in position $m-n$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$.
-If $\sigma_{k,t-1}<m-n\leq\sigma_{k,t}$ for some $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1, then $u$ is a $(D,k,D-t)$-type monomial.
-Since $0=\sigma_{1,0}<m-n\leq\dim(\mathbb{F}_{q}[x_{1},\ldots,x_{n}]_{D})/(x_{1}^{D},\ldots,x_{n}^{D})_{D}=\sigma_{n-1,D-1}$, then $m-n$ always belong to one of the intervals above.
+$\begin{split}\sigma_{k,t}&=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}+\sum_{j=1}^{t}\binom{n-k-1+j}{j}\\ &=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\dim(\mathbb{F}[x_{i+1},\ldots,x_{n}])_{j}+\sum_{j=1}^{t}\dim(\mathbb{F}[x_{k+1},\ldots,x_{n}])_{j}\end{split}$
+is the number of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]$ different from $x_{1}^{D},\ldots,x_{n}^{D}$ and bigger than or equal to $x_{k}^{D-t}x_{n}^{t}$, the smallest (D,k,Dt)(D,k,D-t)-type monomial.
+In other words, the monomial $x_{k}^{D-t}x_{n}^{t}$ is in position $\sigma_{k,t}$ in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$.
+Notice moreover that $\sigma_{1,0}=0$ and $\sigma_{k,0}=\sigma_{k-1,D-1}$ for $2\leq k\leq n-1$.
+If $u$ is the smallest monomial in $\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle$, then $u$ is in position mnm-n in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$.
+If $\sigma_{k,t-1}<m-n\leq\sigma_{k,t}$ for some $1\leq k\leq n-1$ and $1\leq t\leq D-1$, then $u$ is a (D,k,Dt)(D,k,D-t)-type monomial.
+Since $0=\sigma_{1,0}<m-n\leq\dim(\mathbb{F}[x_{1},\ldots,x_{n}]_{D})/(x_{1}^{D},\ldots,x_{n}^{D})_{D}=\sigma_{n-1,D-1}$, then mnm-n always belong to one of the intervals above.
 The ideal $L$ is generated by the degree $D$ monomials which are greater than or equal to $u$, unless the monomial following $u$ in lexicographic decreasing order is a pure power.
-In that case, $u=x_{k}x_{n}^{D-1}$ for some $11\leq k\leq n-$ 1, i.e.
+In that case, $u=x_{k}x_{n}^{D-1}$ for some $1\leq k\leq n-1$, i.e.
 $m-n=\sigma_{k,D-1}$, and the smallest degree $D$ monomial in $L$ is $x_{k+1}^{D}$.
@@
 where the inequality follows from Conjecture 20. ∎
+We stress that the upper bound from Theorem 23 is sharp for all values of $m,n,D$.
+In fact, it is met by any system $\mathcal{F}$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ is a $(D,\ldots,D;D)$-LPP ideal.
+In the next example we show how to compute the bound from Theorem 23 for concrete choices of the parameters.
 Example 24.
-In this example we show how to compute the bound from Theorem 22 for concrete choices of the parameters.
-Let $n=6$ and $D=3$, so =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{6}]_{3}$.
-Then $1\leq k\leq 5$ and $0\leq t\leq 2$.
-The values of $\sigma_{k,t}$ are
+Let $n=6$ and $D=3$, so $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{6}]_{3}$. Then $1\leq k\leq 5$ and $0\leq t\leq 2$. The values of $\sigma_{k,t}$ are
 $\sigma_{k,t}$ $t=0$ $t=1$ $t=2$
@@
 $k=5$ 48 49 50
-If $m=12$, then $m-n=6$ and $\sigma_{1,1}<6\leq\sigma_{1,2}$. Hence $k=1$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 11.$
-If $m=42$, then $m-n=36$ and $\sigma_{3,0}<36\leq\sigma_{3,1}$. Hence $k=3$, $t=1$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 8.$
-If $m=54$, then $m-n=48$ and $\sigma_{4,1}<48\leq\sigma_{4,2}$. Hence $k=4$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 5.$
+If $m=12$, then mn=6m-n=6 and $\sigma_{1,1}<6\leq\sigma_{1,2}$. Hence $k=1$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 11.$
+If $m=42$, then mn=36m-n=36 and $\sigma_{3,0}<36\leq\sigma_{3,1}$. Hence $k=3$, $t=1$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 8.$
+If $m=54$, then mn=48m-n=48 and $\sigma_{4,1}<48\leq\sigma_{4,2}$. Hence $k=4$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 5.$
 Remark 25.
 The upper bound of Theorem 23 is decreasing as a function of $m$, as one would expect.
-In particular, as $m-n$ passes from an interval $(\sigma_{k,t-1},\sigma_{k,t}]$ to the next, the upper bound decreases by one.
+In particular, as mnm-n passes from an interval $(\sigma_{k,t-1},\sigma_{k,t}]$ to the next, the upper bound decreases by one.
 The largest value of the bound is obtained in the case $m=n$, which corresponds to $\mathcal{F}^{\operatorname{\mathrm{top}}}$ being a regular sequence.
-In this case, the value for the bound is well-know and is $n(D-1)+1$.
+In this case, the value for the bound is well-known and is n(D1)+1n(D-1)+1.
 The smallest value for the bound is obtained in the case $m-n=\sigma_{n-1,D-1}$, which corresponds to $\langle\mathcal{F}^{\operatorname{\mathrm{top}}}\rangle+\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle=R_{D}$.
-In this case, the value for the bound is easily seen to be $D$.
-Remark 25.
-The upper bound produced in Theorem 22 is sharp for all values of $m,n,D$.
-In fact, it is met by any system $\mathcal{F}$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ is a $(D,\ldots,D;D)$-LPP ideal.
-Combining Theorem 22 and Theorem 4, we obtain the following.
+In this case, the value of the bound is $D$.
+In order to provide a concrete example, we apply Theorem 23 to the Cubic Simple Matrix encryption scheme proposed in [26].
+This provides the only proven bound on the degree of regularity of this system, to the best of our knowledge.
+Example 26.
+The Cubic Simple Matrix encryption scheme consists of $m=2n$ cubic equations in $n$ variables, with $n$ a square.
+To compute the bound of Theorem 23 for this system, we need to find the unique $1\leq k\leq n-1$ and $t\in\{1,2\}$ such that $m-n=n\in(\sigma_{k,t-1},\sigma_{k,t}]$.
+Since $\sigma_{1,1}=n-1$ and $\sigma_{1,2}=n-1+{n\choose 2}$, we find $k=1$ and $t=2$.
+Hence, the Cubic Simple Matrix encryption scheme has $d_{\operatorname{reg}}\leq 2n-1$.
+We are not aware of any other proven bound on the degree of regularity of this system.
+However, [26] contains a heuristic analysis for $n\gg 0$ and experimental results for small $n$.
+In [26, Section 4.2] the authors perform a heuristic analysis and conclude that, for $n$ sufficiently large, the complexity of solving the cubic system using the XL Algorithm should be at least that of solving a random quadratic system of $n$ equations in $n$ variables.
+The degree of regularity of the latter is $n+1$, provided that the top degree part of the equations is a regular sequence.
+In the same section, they carry on numerical experiments for $n=4,9,16$, from which they find degree of regularity $5,6,7$, respectively.
+Combining Theorem 23 and Theorem 4, we obtain the following. A related result concerning the solving degree of a standard algorithm will be proven after Theorem 33.
 Corollary 27.
-Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials.
-If $m=n$, then let $k=0$, $t=D-1$.
-Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a system of polynomials such that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$.
+Assume that $d_{\operatorname{reg}}(\mathcal{F})<+\infty$.
+If $m=n$, then let $k=0$, t=D1t=D-1.
+Else, let $1\leq k\leq n-1$ and $1\leq t\leq D-1$ be such that mnm-n belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$.
 If Conjecture 20 holds, then
@@
 The upper bound found in Theorem 23 for the degree of regularity of $\mathcal{F}$ is bigger than or equal to $D$ for all $k$ and $t$. The thesis then follows from Theorem 4. ∎
-We now wish to apply Theorem 22 to a system which contains the field equations.
 Remark 28.
-After reducing the equations of $\mathcal{F}$ modulo the field equations, they have degree at most $q-1$ in each variable, hence total degree at most $n(q-1)$.
-Therefore, when adding the field equations to a polynomial system of degree $D$, we may always assume that
+While the estimates of Theorem 23 and Corollary 27 hold for every $D$, they are most relevant for $D<q$.
+In fact, if $D\geq q$, the bound on the solving degree of a mutant system obtained using Theorem 4 does not increase if one adds the field equations to the system.
+Moreover, one always has $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq d_{\operatorname{reg}}(\mathcal{F})$.
+Therefore, if $D\geq q$, one should add the field equations to the system before solving it.
+We now bound the degree of regularity and the solving degree of a system of the form $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$, where $\mathcal{F}$ is a system of $m$ equations of degree $D$ in $n$ variables.
+This is especially relevant in the case $\mathbb{F}=\mathbb{F}_{q}$, but the results that we will prove hold (and are stated) over an arbitrary field $\mathbb{F}$.
+Because of Remark 28, we focus on the case when $D\geq q$.
+After reducing the equations of $\mathcal{F}$ modulo the field equations, they have degree at most q1q-1 in each variable, hence total degree at most n(q1)n(q-1).
+Therefore, in the sequel we may assume without loss of generality that
 $D\leq n(q-1).$
-Combining Theorem 22 and Theorem 3, one obtains the following. By Remark 27, we may assume without loss of generality that $D\leq n(q-1)$.
-Corollary 28.
-Assume that Conjecture 19 holds.
-Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials, with $D\leq n(q-1)$.
-If $m=n$, then let $k=0$, $t=D-1$.
-Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$.
-If $d_{\operatorname{reg}}(\mathcal{F})\geq q$, then
-$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2\min\{n(q-1),(n-k+1)(D-1)-t\}.$
-Proof.
-First, since $(x_{1}^{q},\ldots,x_{n}^{q})_{n(q-1)+1}=R_{n(q-1)+1}$, one has
-$d_{\operatorname{reg}}(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})\leq n(q-1)+1.$
-If $D<q$, then
-$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{q-1}=(\mathcal{F}^{\operatorname{\mathrm{top}}})_{q-1}\neq R_{q-1},$
-since $d_{\operatorname{reg}}(\mathcal{F})\geq q$ by assumption.
-Therefore, $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq q=\max\{D,q\}$.
-If $q\leq D\leq n(q-1)$, then
-$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1},$
-hence $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq D=\max\{D,q\}$.
-This shows that the assumptions of Theorem 3 are satisfied.
-The thesis now follows by combining Theorem 22, Theorem 3, and the observation that
-$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq d_{\operatorname{reg}}(\mathcal{F}).$
-Remark 29.
-While the estimates of Corollary 26 and of Corollary 28 hold for every $D$, they are most relevant for $D\leq q$.
-In fact, for $D>q$ we obtain tighter upper bounds on the degree of regularity - hence on the solving degree - of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in Theorem 34 by inspecting the degree $D$ part of the system $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}$.
-This corresponds to the fact that, whenever the degree of the equations of the system is larger than or comparable to the field size, it is convenient to add the field equations to the system before computing a Gröbner basis.
-Next we estimate the degree of regularity and the solving degree of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in the case when $D\geq q$.
-By Remark 27, we may assume that $D\leq n(q-1)$.
 First, we observe that one can easily derive a lower bound on the degree of regularity.
 Remark 29.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $q\leq D\leq n(q-1)$.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D\leq n(q-1)$.
 Since $(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1}$, then
@@
 We now want to derive an upper bound. We start by computing the number of linearly independent homogeneous polynomials of degree $D$ as a function of $n,D,q$.
-Remark 31.
+Lemma 30.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{1})=\ldots=\deg(f_{m})=D$.
 Assume that $q\leq D\leq n(q-1)$.
+If $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$, then
+$m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$. (3)
+Conversely, if inequality (3) is satisfied, then the set of systems $\mathcal{F}$ such that that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is open and nonempty.
+Proof.
 A standard Hilbert function computation shows that the number of homogeneous polynomials in $n$ variables of degree $D$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is
 $\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}=\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$.
-It follows that the assumption that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ holds on a dense open set, whenever
-$m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}.$ (2)
-If inequality (2) is not satisfied, then $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ cannot be linearly independent.
-The simplest case to treat is that of very overdetermined systems, more specifically the case when $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are too many to be linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$.
-This happens when $m$ is larger than the bound from Remark 31.
-The next proposition shows that, in such a situation, the degree of regularity is equal to the degree of the equations of the system.
+Therefore, if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent, then inequality (3) is satisfied.
+The set of systems $\mathcal{F}$ as in the statement is always open, since linear dependence is a rank condition, hence a closed condition.
+Suppose then that inequality (3) is satisfied.
+Then there exist $g_{1},\ldots,g_{m}\in(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ linearly independent, so the set of systems $\mathcal{F}$ in the statement is nonempty.
+∎
+The simplest case is that of very overdetermined systems, more specifically the case when $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are too many to be linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$.
+This happens when $m$ is larger than the bound from Lemma 30.
+It is easy to show that, in such a situation, the degree of regularity is equal to the degree of the equations of the system.
 Proposition 31.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, where $q\leq D\leq n(q-1)$.
-If $m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}$, then there is a dense open set $\mathcal{W}$ such that, if $\mathcal{F}\in\mathcal{W}$, then
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, where $q\leq D\leq n(q-1)$.
+If $m\geq\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}$, then there is a dense open set $\mathcal{W}$ such that, if $\mathcal{F}\in\mathcal{W}$, then
 $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})=D.$
@@
 Since by assumption
-$m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}=\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D},$
+$m\geq\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}=\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D},$
 then there is an open set $\mathcal{W}$ of $m$-tuples of polynomials of degree $D$ such that
@@
 The rest of the statement now follows from Theorem 3 and Theorem 4. ∎
-The next theorem yields an upper bound on the degree of regularity of an algebraically random system of equations of degree larger than the field size to which we add the field equations.
-For such a system, the bound is tighter than the one from Corollary 28.
+The next theorem yields an upper bound on the degree of regularity of a system of equations of degree $D$ larger than the field size, to which we add the field equations.
 We start with a preparatory lemma, whose proof follows directly from the definition.
@@
 In particular, if $u,v,w$ are monomials such that $u\geq v\geq w$ and $u$ and $w$ have the same type, then $v$ also has the same type as $u$ and $w$.
+The next theorem applies, e.g., to a system of equations of the same degree $D\geq 2$ over $\mathbb{F}_{2}$.
+Notice that, by Lemma 30, the assumption that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$ implies that $m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$.
 Theorem 33.
 Assume that Conjecture 20 holds.
-Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, with $q\leq D\leq n(q-1)$.
-Assume that $m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$ and that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$.
-Let $11\leq t\leq q-$ 1 and $11\leq k\leq n-$ 1 be such that $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
-$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, with $q\leq D\leq n(q-1)$.
+Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$.
+Let $1\leq t\leq q-1$ and $1\leq k\leq n-1$ be such that $m$ belongs to the interval $(\varsigma_{k,t-1},\varsigma_{k,t}]$, where
+$\varsigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
 and
@@
 Let
-$B=q-t+(n-k)(q-1).$
-If $m=\sigma_{k,t}$ and either $t\neq q-1$ and $D\geq 2q-t-1$, or $t=q-1$ and $D\geq 2q-1$, then
+B=qt+(nk)(q1).B=q-t+(n-k)(q-1).
+If $m=\varsigma_{k,t}$ and either $t\neq q-1$ and $D\geq 2q-t-1$, or t=q1t=q-1 and $D\geq 2q-1$, then
 $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B-1,$
@@
 and
-$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B+1.$
+$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq B.$
 In any other case
@@
 Proof.
-We start by proving that $\eta_{k,t}$ is the number of $(D,k,q-t)$-type monomials in $R$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$, that is
-$\eta_{k,t}=\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}.$ (3)
-Since $x_{k+1}^{q},\ldots,x_{n}^{q}$ is a regular sequence in $\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]$, a standard computation involving Hilbert series yields the explicit formula
-$\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1},$
+The idea of the proof is similar to that of Theorem 23.
+In order to apply Proposition 21, we need to determine the value of $k$ and $t_{k}$ for the monomial in position $m$ in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$.
+As we will show, these values are the unique $k$ and $t_{k}$ such that $m$ belongs to the interval $(\varsigma_{k,q-t_{k}-1},\varsigma_{k,q-t_{k}}]$.
+We start by proving that $\eta_{k,t}$ is the number of (D,k,qt)(D,k,q-t)-type monomials in $R$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$, that is
+$\eta_{k,t}=\dim\left[\frac{\mathbb{F}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}.$ (4)
+Since $x_{k+1}^{q},\ldots,x_{n}^{q}$ is a regular sequence in $\mathbb{F}[x_{k+1},\ldots,x_{n}]$, a standard computation involving Hilbert series yields the explicit formula
+$\dim\left[\frac{\mathbb{F}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1}$,
 where we notice that $\eta_{k,t}\neq 0$ only if $D+t-q\leq(n-k)(q-1)$. This establishes the equality in (4). Similarly to the proof of Theorem 23, we notice that
-$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
-is the number of monomials of degree $D$ which do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$ and are greater than or equal to $x_{k}^{q-t}x_{n}^{D-q+t}$, the smallest $(D,k,q-t)$-type monomial.
-Moreover, $\sigma_{k-1,q-1}=\sigma_{k,0}$ for $12\leq k\leq n-$ 1.
-Let $I\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\}$.
+$\varsigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
+is the number of monomials of degree $D$ which do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$ and are greater than or equal to $x_{k}^{q-t}x_{n}^{D-q+t}$, the smallest (D,k,qt)(D,k,q-t)-type monomial.
+Moreover, $\varsigma_{k-1,q-1}=\varsigma_{k,0}$ for $2\leq k\leq n-1$.
+Let $I\subseteq\mathbb{F}[x_{1},\ldots,x_{n}]$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\}$.
 Let $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$ be the $(q,\ldots,q;D)$-LPP ideal such that
@@
 Hence $\mathcal{L}$ is minimally generated by $x_{1}^{q},\ldots,x_{n}^{q}$ and $m$ monomials of degree $D$ that do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$.
-Recall that, by assumption, $L$ is the largest lex-segment ideal generated in degree $D$ such that $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$, i.e., $L_{D}\supseteq(x_{1}^{q},\ldots,x_{n}^{q})_{D}$.
+Recall that, by assumption, $L$ is the largest lex-segment ideal generated in degree $D$ such that $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$.
 Let $u$ be the smallest degree $D$ monomial in $\mathcal{L}/(x_{1}^{q},\ldots,x_{n}^{q})$.
-Notice that $u$ is a $(D,k,q-t)$-type monomial, since $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$.
+Notice that $u$ is a (D,k,qt)(D,k,q-t)-type monomial, since $m$ belongs to the interval $(\varsigma_{k,t-1},\varsigma_{k,t}]$.
 Let $v$ be the smallest degree $D$ monomial in $L$.
-If $u$ is the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$, then $u=x_{k}^{q-t}x_{k+1}^{q-1}\cdots x_{n}^{q-1}$ and $D=q-t+(n-k)(q-1)$.
+If $u$ is the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$, then $u=x_{k}^{q-t}x_{k+1}^{q-1}\cdots x_{n}^{q-1}$ and D=qt+(nk)(q1)D=q-t+(n-k)(q-1).
 Moreover, $\mathcal{L}_{D}=R_{D}$ and $v=x_{n}^{D}$ is a $(D,n,D)$-type monomial.
@@
 Then $u\geq v\geq w$ and $v$ is the degree $D$ monomial next to $w$ in increasing lexicographic order.
-If $m\neq\sigma_{k,t}$, then $w$ has type $(D,k,q-t)$, hence so does $v$ by Lemma 33.
-Suppose therefore that $m=\sigma_{k,t}$.
-Write $D-q+t=(n-\ell)(q-1)+r$, where $10\leq r<q-$ 1.
-Notice that $D-q+t<(n-k)(q-1)$, since $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$.
+If $m\neq\varsigma_{k,t}$, then $w$ has type (D,k,qt)(D,k,q-t), hence so does $v$ by Lemma 32.
+Suppose therefore that $m=\varsigma_{k,t}$.
+Write $D-q+t=(n-\ell)(q-1)+r$, where $0\leq r<q-1$.
+Notice that Dq+t<(nk)(q1)D-q+t<(n-k)(q-1), since $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$.
 Then $\ell>k$.
 In this situation, $u=x_{k}^{q-t}x_{\ell}^{r}x_{\ell+1}^{q-1}\cdots x_{n}^{q-1}$ and $w=x_{k}^{q-t-1}x_{k+1}^{q-1}\cdots x_{k+n-\ell}^{q-1}x_{k+n-\ell+1}^{r+1}$.
-Notice that $w$ is a $(D,k,q-t-1)$-type monomial if $t\neq q-1$ and a $(D,k+1,q-1)$-type monomial if $t=q-1$.
+Notice that $w$ is a (D,k,qt1)(D,k,q-t-1)-type monomial if $t\neq q-1$ and a (D,k+1,q1)(D,k+1,q-1)-type monomial if t=q1t=q-1.
 If $w$ is not the smallest degree $D$ monomial of its type in $R_{D}$, then $v$ has the same type as $w$.
-If $w$ is the smallest degree $D$ monomial of its type in $R_{D}$, then $w=x_{k}^{q-t-1}x_{k+1}^{D-q+t+1}$ in the case $t\neq q-1$ and $w=x_{k+1}^{q-1}x_{k+2}^{D-q+1}$ in the case $t=q-1$.
-Since $w\not\in(x_{1}^{q},\ldots,x_{n}^{q})$, this is only possible if $D\leq 2q-t-2$ for $t\neq q-1$ and $D\leq 2q-2$ for $t=q-1$.
-In this situation, $v$ has type $(D,k,q-t)$ in the case $t\neq q-1$ and type $(D,k+1,q)$ in the case $t=q-1$.
+If $w$ is the smallest degree $D$ monomial of its type in $R_{D}$, then $w=x_{k}^{q-t-1}x_{k+1}^{D-q+t+1}$ in the case $t\neq q-1$ and $w=x_{k+1}^{q-1}x_{k+2}^{D-q+1}$ in the case t=q1t=q-1.
+Since $w\not\in(x_{1}^{q},\ldots,x_{n}^{q})$, this is only possible if $D\leq 2q-t-2$ for $t\neq q-1$ and $D\leq 2q-2$ for t=q1t=q-1.
+In this situation, $v$ has type (D,k,qt)(D,k,q-t) in the case $t\neq q-1$ and type $(D,k+1,q)$ in the case t=q1t=q-1.
 If Conjecture 20 holds, then
@@
 Moreover, by Proposition 21
-$\operatorname{reg}(\mathcal{L})=\left\{\begin{array}[]{ll}q-t-1+(n-k)(q-1)&\mbox{if }m=\sigma_{k,t}\mbox{ and }\\ &\mbox{either }t\neq q-1\mbox{ and }D\geq 2q-t-1,\\ &\mbox{or }t=q-1\mbox{ and }D\geq 2q-1,\\ \\ q-t+(n-k)(q-1)&\mbox{else.}\end{array}\right.$ (5)
+[…]
 The bound on the degree or regularity now follows from (5) and (6). The bounds on the solving degree follow from the bound on the degree of regularity, Theorem 3, and Theorem 4. ∎
-5 Applications to the study of GeMSS and Rainbow
-GeMSS and Rainbow were the only multivariate schemes in Round 3 of the NIST Post-Quantum Cryptography Standardization process.
-They are based on modifications of HFE (Hidden Field Equation) and UOV (Unbalanced Oil and Vinegar), respectively.
-They were subsequently revealed to be insecure and susceptible to a MinRank Attack, see e.g.
-[3, 8, 9, 1], therefore they were excluded from the NIST competition.
-In this section, we demonstrate how to use our results to show that these systems are far from being algebraically random.
-We do so by computing their degree of regularity or solving degree for small instances and comparing it with the upper bounds that we obtained in Theorem 22 and Theorem 34.
-The degree of regularity and solving degree of the systems associated to GeMSS and Rainbow are much smaller than the corresponding invariants for an algebraically random system with the same parameters, which reveals the presence of a hidden structure that may be used to mount an ad-hoc attack, as it was done in practice with the MinRank Attacks mentioned above.
-We used Magma to compute the solving degree and Singular to compute the degree of regularity.
-The values that we obtain (and that we indicate in the tables below) are almost always the same for systems with the same parameters.
-For each choice of the parameters in the table, we produce ten instances of the public key $\mathcal{PK}=\{p_{1},\ldots,p_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of the chosen scheme.
-For each one of them, we choose a random vector $s=(s_{1},\ldots,s_{m})\in\mathbb{F}_{q}^{m}$ as a signature (in case the chosen vector is not a valid signature, we replace it with another randomly chosen vector).
-In order to forge the signature, an attacker may want to solve the system $\mathcal{PK}s=\{p_{1}-s_{1},\ldots,p_{m}-s_{m}\}$.
-We make the system $\mathcal{PK}s$ square by assigning random values to the last $n-m$ variables.
-This yields a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$.
-Since in GeMSS we work over $\mathbb{F}_{2}$, we add to the system $\mathcal{F}$ the field equations =$\mathcal{E}=\{x_{1}^{2}+x_{1},\ldots,x_{m}^{2}+x_{m}\}$.
-In the next table we compare the experimental results we obtained for GeMSS with the bounds from Theorem 34.
-The experiments show that both the solving degree and the degree of regularity of $\mathcal{F}$ can be more than twice the solving degree of $\mathcal{F}\cup\mathcal{E}$.
-This confirms the intuition that adding the field equations is a good strategy in order to solve the system $\mathcal{F}$ over $\mathbb{F}_{2}$.
-Unfortunately we were able to compute the degree of regularity of $\mathcal{F}\cup\mathcal{E}$ only for small values of the parameters.
-In the next table the first three columns contain the parameters of the cryptosystem, and the fourth the number of polynomials and variables that appear in $\mathcal{F}$.
-The columns labelled '$d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular.
-The columns labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' and '$\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' are the bounds given by the Theorem 34 for the chosen parameters and a standard Gröbner basis algorithm.
-n, D a v m $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
-12, 4 1 1 11 5 10 3 18
-8, 9 1 1 7 3 6 3 10
-8, 9 1 2 7 3 6 3 10
-8, 9 2 1 6 3 5 3 8
-8, 9 2 2 6 3 5 3 8
-24, 4 1 1 23 22 4 42
-24, 4 1 2 23 22 4 42
-24, 4 1 3 23 22 4 42
-24, 4 2 1 22 21 4 40
-24, 4 2 2 22 21 4 40
-24, 4 3 1 21 20 4 38
-While GeMSS is random according to Definition 12, the experimental results make it clear that both the degree or regularity and the solving degree of GeMSS are far from the largest values that one can find for a system of those parameters according to Theorem 34.
-This indicates that, for the same parameters, one should be able to find systems for which the complexity of computing a Gröbner basis is much larger.
-More importantly, it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the MinRank attacks mentioned in the opening paragraph).
-For Rainbow, we choose to work over $\mathbb{F}_{4}$ and $\mathbb{F}_{9}$.
-Since $\mathcal{F}$ is a square system, $\mathcal{F}$ is algebraically random if and only if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence.
-This turns out to be the case in most of the examples that we computed and in that case
-$d_{\operatorname{reg}}(\mathcal{F})=m+1$ (6)
-by Theorem 22. This is confirmed by our computations.
-Since the systems coming from this scheme are quadratic, adding the field equations may increase the solving degree of the system.
-However, for the small values of $q$ that we tried in our experiments, we find that in all cases but one the solving degree decreases when adding the field equations.
-This makes sense, as the degree $q$ of the equations that we add is never larger than the solving degree of the system to which we add them.
-In the next table we summarize the results that we obtained in our computational experiments.
-Since in our experiments the solving degree of $\mathcal{F}\cup\mathcal{E}$ is almost always smaller than that of $\mathcal{F}$, in our examples $\mathcal{F}\cup\mathcal{E}$ is the relevant system to consider, that is, the system that one wants to try to solve.
-Therefore, we consider the degree of regularity and solving degree of $\mathcal{F}\cup\mathcal{E}$.
-In our table, we compare the degree of regularity and the solving degree of $\mathcal{F}\cup\mathcal{E}$ with the upper bounds from Theorem 22 and Corollary 28.
-We use Corollary 28 as the algorithm implemented in Magma is a standard one.
-The first three columns contain the chosen values for the parameters and the number of polynomials and variables that appear in $\mathcal{F}$.
-The columns labelled '$d_{\operatorname{reg}}(\mathcal{F})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular.
-The column labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contains the bounds from Corollary 28.
-We do not include the values of the degree of regularity and solving degree of $\mathcal{F}$ in the table, as experimentally we find that the system $\mathcal{F}\cup\mathcal{E}$ can always be solved more efficiently than the system $\mathcal{F}$.
-$q$ $[v_{1},o_{1},o_{2}]$ $m$ $d_{\operatorname{reg}}(\mathcal{F})$ $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
-4 $[3,2,2]$ 4 4/5 4 4 8
-4 $[3,3,3]$ 6 6/7 5 5 12
-4 $[7,5,5]$ 10 10/11 6 6 20
-9 $[3,2,2]$ 4 5 5 9
-9 $[7,5,5]$ 10 11 9/10 10 20
-Notice that Corollary 28 does not apply to the case $q=9$ and $[v_{1},o_{1},o_{2}]=[3,2,2]$, since $d_{\operatorname{reg}}(\mathcal{F})=5$.
-For these parameters, the bound from Corollary 28 would yield $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})\leq 8$.
-However, the bound does not hold in this case, as our experiments show.
-As for GeMSS, we observe that the values that we computed for the solving degree of $\mathcal{F}\cup\mathcal{E}$ are far from the upper bounds predicted by Corollary 28.
-We conclude that, also in this case, one expects to find systems with the same parameters as these instances of Rainbow and for which the complexity of computing a Gröbner basis is larger.
-More importantly, once again it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the attacks that we mentioned at the start of the section).
+We state as a corollary the case of a system of binary equations. This is a case of special interest and has a simpler formulation than the general case.
+Corollary 34.
+Assume that Conjecture 20 holds.
+Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a system of degree $D$ polynomials in $n$ variables over $\mathbb{F}_{2}$, with $2\leq D\leq n$.
+Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $x_{1}^{2},\ldots,x_{n}^{2}$.
+Let $1\leq k\leq n-1$ be such that $m$ belongs to the interval $(\varsigma_{k-1},\varsigma_{k}]$, where
+$\varsigma_{k}=\sum_{i=1}^{k}{n-i\choose D-1}$.
+If $m=\varsigma_{k}$ and and $D\geq 3$, then
+$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k,$
+$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq 2(n-k-1),$
+and
+$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k+1.$
+Else
+$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k+1,$
+$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq 2(n-k),$
+and
+$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq n-k+2.$
+Corollary 33 applies to, e.g., the Unbalanced Oil and Vinegar scheme (UOV) from [32, Section 2], which over $\mathbb{F}_{2}$ is a system of $m$ quadratic equations in $n>m$ variables, to which one adds the $n$ quadratic field equations.
+It also applies to the Oil and Vinegar scheme of degree three from [32, Section 9] and to the Unbalanced Oil, Vinegar and Salt scheme from [32, Section 11].
+Over $\mathbb{F}_{2}$, the first is a system of $m$ cubics in $n>m$ variables, to which one adds the $n$ quadratic field equations.
+The second is a system of $m$ quadratic equations with the Oil and Vinegar structure in $n>m$ variables, to which one adds $r$ random quadratic equations and the $n$ quadratic field equations.
+Example 35.
+Let $\mathcal{F}$ be an UOV system in degree $2\leq D\leq n$ over $\mathbb{F}_{2}$.
+For a UOV system one always has $m<n$, since nmn-m is the number of vinegar variables.
+If $D=n$, then $m=\varsigma_{1}=1$.
+Else, $\varsigma_{1}={n-1\choose D-1}\geq n-1$.
+In both cases, one has $m\in(0,\varsigma_{1}]$ and $k=1$.
+Therefore, by Corollary 34
+$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n$
+and if either $D=n$ and $m=1$, or $D\geq 3$ and m=n1m=n-1, then
+$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-1.$
+One can also approach a UOV system by assigning random values to nmn-m variables and trying to solve for the rest.
+This yields a system of $m$ equations in $m$ variables, to which one can still add the field equations.
+We compute the bound for a quadratic UOV system.
+Example 36.
+Let $\mathcal{F}$ be an UOV system in degree $D=2$ over $\mathbb{F}_{2}$.
+Assign random values to nmn-m variables to obtain a system of $m$ quadratic equations in $m$ variables.
+If $m>2$, then $m\leq\dim_{\mathbb{F}_{2}}\mathbb{F}_{2}[x_{1},\ldots,x_{m}]_{2}-\dim_{\mathbb{F}_{2}}\langle x_{1}^{2},\ldots,x_{m}^{2}\rangle$, so one expects that the specializations of $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ remain linearly independent modulo $\langle x_{1}^{2},\ldots,x_{m}^{2}\rangle$.
+Since $\varsigma_{1}=m-1$ and $\varsigma_{2}=2m-3$, then $m\in(\varsigma_{1},\varsigma_{2}]$ and $k=2$.
+Therefore, by Corollary 34 the degree of regularity is bounded from above by m1m-1.
+In Theorem 23 and Theorem 33 we consider a system $\mathcal{F}$ of equations of the same degree $D$ and provided upper bounds on the degree of regularity of $\mathcal{F}$ under the assumption that $\mathcal{F}$ is regular in degree $D$, and of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ if $D\geq q$.
+Since our results also apply to semiregular systems, the bounds that we obtain are necessarily larger than the value of the degree of regularity of a semiregular system of equations of the same degree.
+A common approach in the cryptoanalysis of a multivariate scheme is to assume that the corresponding system is semiregular.
+Under such an assumption, one can compute the value of the degree of regularity of the system.
+We stress that, even though the bounds that we obtain are less tight than those that one obtains by assuming that a system is semiregular, they have the advantage of being provable, since they just require that the degree of regularity is finite and do not rely on the heuristic assumption that the system is semiregular.